### v1

# On the distance between mean and geometric median in high dimensions

Abstract
The geometric median, a notion of “center” for multivariate distributions, has gained recent attention in robust statistics and machine learning. Although conceptually distinct from the mean (i.e., expectation), we demonstrate that both are very close in high dimensions when the dependence between the distribution components is suitably controlled. Concretely, we find an upper bound on the distance that vanishes with the dimension asymptotically, and derive a rate-matching first order expansion of the geometric median components. Simulations illustrate and confirm our results.
keywords:
spatial median, robust mean estimation
1 Introduction
The geometric median provides a robust alternative to the sample mean, mitigaging the impact of outliers. The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which in practice often is the learning target [^1] For special distributions like the Gaussian, geometric median and mean can coincide. . A method to balance robustness and bias is the median-of-means [Lugosi2019, Minsker2015] or “bucketing” approach [Guerraoui2024].
In this paper, we take a step back and quantify the actual bias of the geometric median, focusing on high-dimensional distributions. We prove, when each distribution component only depends on a fixed number of neighboring components, that the maximum-norm distance between geometric median and mean vector is of order $\mathcal{O}(\tfrac{1}{p})$, where $p$ denotes the dimension. Furthermore, we show that this rate is sharp in general.
The results are discussed at the end of the paper. Proofs are deferred to the appendix.
2 Theoretical results
We consider random vectors $\mathbf{X}:=(X_{1},\ldots,X_{p})\in\mathbb{R}^{p}$ for $p\in\mathbb{N}$, where the sequence of component random variables $(X_{i})_{i\in\mathbb{N}}$ is from the class $\mathcal{D}\equiv\mathcal{D}(M,q,c,\sigma_{\min},C)$ introduced below. A geometric median of $\mathbf{X}$ is any vector $\mathbf{m}\in\mathbb{R}^{p}$ minimizing the expectation $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ for the Euclidean norm $\|\cdot\|$. By properties of the class $\mathcal{D}$, there exists a unique geometric median (Lemma A.1), which we denote by $\mathbf{m}_{p}$. We then study $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|$, where $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$.
Definition 2.1.
Let $M\geq 0$ be an integer, and let $c,\sigma_{\min}>0$, $q\geq 2$ and $C\geq 1$ be reals. The class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ is comprised of all sequences of absolutely continuous r.v. $(X_{i})_{i\in\mathbb{N}}$ that are $M$-dependent, with $\operatorname{E}\left[\,|X_{i}|^{q}\,\right]\leq C$, $\operatorname{E}\left[\,|X_{i}-\mu_{i}|^{q}\,\right]\leq C$, and $\operatorname{sd}\left[\,X_{i}\,\right]\geq\sigma_{\min}>0$, and for which the probability density of $X_{i}$ is bounded by $c<\infty$ for all $i\in\mathbb{N}$.
Recall that $M$-dependence requires independence of $(X_{i})_{i\leq k}$ and $(X_{i})_{i\geq k+M+1}$ for all $k\in\mathbb{N}$. Note that $|\mu_{i}|\leq C$ and $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ since $C\geq 1$.
We build on the following central identity, derived from $\nabla_{\mathbf{m}}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]|_{\mathbf{m}=\mathbf{m}_{p}}=0$ by interchanging derivative with expectation, and then adding $\pm\boldsymbol{\mu}$:
$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right],$ (1)
where all expectations exist when $p$ is large enough (Lemma A.2).
The intuition is now that, by $M$-dependence, $X_{i}-\mu_{i}$ and $\|\mathbf{X}-\mathbf{m}_{p}\|$ are nearly independent for $p$ large, and thus $m_{p,i}-\mu_{i}\approx\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\cdot 1=0$ by (1). Formalizing this approach requires controlling $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{-1}\,\right]$, or slight variations involving $\check{\mathbf{X}}$ introduced below, uniformly in the dimension $p$. To this end, we combine the two lemmas, where the first is an initial bound.
Lemma 2.2.
When $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$, then $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\sqrt{Cp}$.
Lemma 2.3.
Let $(Y_{i})_{i\in\mathbb{N}}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$ and $\mathbf{v}_{p}\in\mathbb{R}^{p}$ with $\|\mathbf{v}_{p}\|^{2}\leq C_{2}p$ for $p\in\mathbb{N}$ and some $C_{2}<\infty$. With $\mathbf{Y}:=(Y_{1},\ldots,Y_{p})$ we have
$\sup_{p\geq\lceil 4k\rceil(M+1)+1}\operatorname{E}\left[\,\left(\frac{\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}\,\right]}\right)^{-k}\,\right]=\beta(k,C_{2},\mathcal{D})<\infty,$ (2)
for any $k>0$ and constant $\beta$ depending only on $k,C_{2}$ and the $\mathcal{D}$-parameters.
Both Lemmas together with the Cauchy-Schwarz inequality applied to the second expectation in (1) lead to the following intermediate bound.
Lemma 2.4.
If $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, then for a constant $B_{1}$ only depending on the $\mathcal{D}$-parameters, $\sup_{p\in\mathbb{N};\;1\leq i\leq p}|\mu_{i}-m_{p,i}|\leq B_{1}$.
We can now derive the main results, which refine the Cauchy-Schwarz bound in the proof of Lemma 2.4. More precisely, we show that $\operatorname{E}\left[\,(X_{i}-\mu_{i})/\|\mathbf{X}-\mathbf{m}_{p}\|\,\right]$ is close to zero by inserting $0=\operatorname{E}\left[\,(X_{i}-\mu)/\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\,\right]$, where $\check{\mathbf{X}}$ is missing all components with indices in $i-M,\ldots,i+M$. Formally, using the exclusion set $\mathcal{E}(i):=\mathbb{N}\cap\{i-M,\ldots,i+M\}$, we define $\check{\mathbf{v}}:=(v_{i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}(i)}\in\mathbb{R}^{|\{1,\ldots,p\}\setminus\mathcal{E}(i)|}$ for any vector $\mathbf{v}=(v_{i})_{i=1}^{p}\in\mathbb{R}^{p}$. Thus, subtracting zero from the right hand side of (1), a bit of algebra yields that for $p\geq i+M$,
$m_{p,i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right].$ (3)
The first expectation is $\mathcal{O}(\sqrt{p})$ while the second is $\mathcal{O}(\tfrac{1}{p^{3/2}})$, leading to
Theorem 2.5.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$. Then, with $B_{3}$ depending only on the $\mathcal{D}$-parameters, $|m_{p,i}-\mu_{i}|\leq\frac{B_{3}}{p}$ for all $1\leq i\leq p\in\mathbb{N}$.
In particular, Theorem 2.5 implies $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\tfrac{B_{3}}{\sqrt{p}}$.
By $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\to 0$, we can identify the limit in probability of the (suitably scaled) expectation arguments in (3). A uniform integrability argument (requiring $q>3$) transfers the convergence to the expectations themselves and yields
Theorem 2.6.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>3$. Fix $i\in\mathbb{N}$. If $\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\to\bar{\sigma}^{2}\in\mathbb{R}$, then
$m_{p,i}=\mu_{i}-\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right]+o\left(\frac{1}{p}\right).$ (4)
Note that while $\sigma_{\min}^{2}\leq\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ always has a convergent subsequence, there may be multiple accumulation points.
3 Simulations
We simulate $\hat{\mathbf{m}}_{p}$ for a range of distributions to validate Theorems 2.5 and 2.6. The simulation details are discussed at the end of this section.
Figure 1: (Left) Euclidean distance $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$. (Right) Residuals from Theorem 2.6 for $i=1$; see equation (5). Black reference lines are $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ (left) and $\mathcal{O}(\tfrac{1}{p})$ (right).
We plot $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$, for which Theorem 2.5 predicts $\mathcal{O}(\tfrac{1}{\sqrt{p}})$, on the left of Figure 1. On the right, we plot the residual
$\left|\hat{m}_{p,1}-\mu_{1}+\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(1)}\operatorname{E}\left[\,(X_{1}-\mu_{1})(X_{j}-\mu_{j})^{2}\,\right]\right|,$ (5)
for which Theorem 2.6 predicts $o(\tfrac{1}{p})$. We consider four data-generating distributions $(X_{i})_{i\in\mathbb{N}}$: first, independent exponential random variables with rates uniformly sampled from $[0.75,1.25]$, second an MA(2) process $X_{i}=L_{i}+0.2L_{i-1}-0.1L_{i-2}$ with iid skew normal noise $(L_{i})_{i}$, third iid Pareto variables with shape parameter $3.1$ (meaning $\operatorname{E}\left[\,|X_{1}|^{k}\,\right]<\infty$ iff $k<3.1$), and finally independent Paretos with shapes uniformly sampled from $[7,8]$. Table 1 lists linear regression slopes to the simulation curves after $p\geq 100$.
The simulation study clearly supports the theory, although $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ approaches $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ rather slowly for the heavy-tailed Pareto distributions. To see why, note $m_{p,1}-\mu_{1}\approx\tfrac{-19.5}{p}+\tfrac{20.2}{p^{1.03}}$ for Pareto(3.1) from the component expansion plot. Since all geometric median components are equal in the iid setting, $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\approx\sqrt{p}\cdot|\tfrac{-19.5}{p}+\tfrac{20.2}{p^{1.03}}|$, which indeed fits the simulations results (Figure 2). The near cancellation in the absolute value slows the decay for moderate $p$.
[…]
4 Discussion
The bias of the geometric median was recently highlighted in [Minsker2024GeometricMedian], where a constant upper bound on the Euclidean bias for a large class of distributions is derived. We complement this result by showing the Euclidean bias to be $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ for a smaller class.
From an estimation viewpoint, a small distance to the mean is desirable. From a robustness viewpoint it raises concern, as the mean is not robust against, say, contaminating a portion of the observations. A differentiation clarifies this: while $M$-dependence is compatible with certain cellwise contamination where indeed the geometric median is not always robust [RAYMAEKERS2024Cellwise], rowwise contaminated observations are typically not $M$-dependent and the geometric median is robust here [Minsker2015, schwank2025robustscorematching].
There is empirical evidence that the geometric median may similarly be close to the mean for sparsely interacting random fields [schwank2025robustscorematching].
Our results link to the classic paper by [Brown1983] showing that (sample) geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
Finally, we remark that the expansion in Theorem 2.6 always contains the term $\operatorname{E}\left[\,(X_{i}-\mu_{i})^{3}\,\right]$, i.e., the negative component skewness. This aligns with statistical intuition for the univariate median [Hippel2005].
Acknowledgements This paper is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research. Further, this work has been funded by the German Federal Ministry of Education and Research and the Bavarian State Ministry for Science and the Arts. The authors of this work take full responsibility for its content.
Appendix A Appendix
A.1 Proofs
Lemma A.1.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ and $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$. Then, when $p\geq M+2$, there exists a unique geometric median of $\mathbf{X}$.
Proof.
For $\mathbf{m},\mathbf{m}^{\prime}\in\mathbb{R}^{p}$, we have that
$|\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]-\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}^{\prime}\|\,\right]|\leq\operatorname{E}\left[\,\|\mathbf{m}-\mathbf{m}^{\prime}\|\,\right]=\|\mathbf{m}-\mathbf{m}^{\prime}\|$
by the inverse triangle inequality, hence $\mathbf{m}\mapsto\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ is continuous. Also, $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ diverges to $\infty$ as $\|\mathbf{m}\|\to\infty$, since
$\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]\geq\operatorname{E}\left[\,|\|\mathbf{X}\|-\|\mathbf{m}\||\,\right]\geq\|\mathbf{m}\|-\operatorname{E}\left[\,\|\mathbf{X}\|\,\right].$
Since continuous functions admit minimizers on compact sets, a geometric median exists.
When $p\geq M+2$, then $X_{1}$ and $X_{M+2}$ are independent, which ensures the distribution of $\mathbf{X}$ does not concentrate on a one-dimensional linear subspace. By [Milasevic1987], the geometric median is unique. ∎
Lemma A.2.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Set $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$, denote $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ and let $\mathbf{m}_{p}$ be the geometric median of $\mathbf{X}$. Then, when $p\geq 2(M+1)+1$,
$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$
Proof.
Let $G(\mathbf{m}):=\|\mathbf{X}-\mathbf{m}\|$ for $\mathbf{m}\in\mathbb{R}^{p}$. We have
$|\partial_{i}G(\mathbf{m})|=\left|\frac{(m_{i}-X_{i})}{\|\mathbf{X}-\mathbf{m}\|}\right|\leq\frac{\|\mathbf{X}-\mathbf{m}\|_{1}}{\|\mathbf{X}-\mathbf{m}\|}\leq\sqrt{p}$
by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [Shorack2017, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
$\nabla\operatorname{E}\left[\,G(\mathbf{m})\,\right]=\operatorname{E}\left[\,\frac{\mathbf{m}-\mathbf{X}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$
Since $\mathbf{m}_{p}$ optimizes $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$, we have
$0=\operatorname{E}\left[\,\frac{\mathbf{X}-\mathbf{m}_{p}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$ (6)
By Lemma 2.2 and 2.3 with $k=\tfrac{1}{2}$ (both proven separately later), we have for $p\geq 2(M+1)+1$ that
$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]\leq\frac{\beta(0.5,C,\mathcal{D})}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\leq\frac{\beta(0.5,C,\mathcal{D})}{\sigma_{\min}\sqrt{p}}<\infty$
by $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\geq\operatorname{E}\left[\,\|\mathbf{X}-\boldsymbol{\mu}\|^{2}\,\right]$ due to MSE optimality of the mean. Therefore, adding $0=\operatorname{E}\left[\,\tfrac{\boldsymbol{\mu}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$ to (6), we find
$\operatorname{E}\left[\,\frac{\mathbf{m}_{p}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]=\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$
Pulling constants out of the expectations and dividing by $\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]>0$ (otherwise $\mathbf{X}=\mathbf{m}_{p}$ almost surely, a contradiction to $\sigma_{\min}>0$), we have shown the claim. ∎
See 2.2
Proof.
By assumption, we have $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ for all $i\in\mathbb{N}$. Using the triangle inequality, optimality of the geometric median, Jensen’s inequality and the variance bound yields
$\|\mathbf{m}_{p}-\boldsymbol{\mu}\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\boldsymbol{\mu}\,\right]\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\mathbf{X}\,\right]\|\leq\operatorname{E}\left[\,\|\mathbf{m}_{p}-\mathbf{X}\|\,\right]\leq\\ \operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|\,\right]\leq\sqrt{\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|^{2}\,\right]}\leq\sqrt{Cp}$
as desired. ∎
See 2.3
Proof.
To simplify notation and stress the additive structure, we denote $S(\mathbf{v}_{p}):=\sum_{i=1}^{p}(Y_{i}-v_{p,i})^{2}=\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}$ and $s_{p}:=\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$ for $p\in\mathbb{N}$. The quantity of interest thus is $\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]$.
Let $k>0$ and $p\geq\lceil 4k\rceil(M+1)+1$. We use $n:=\lfloor\tfrac{p-1}{M+1}\rfloor+1$, and note $n\geq 4k$.
We repeatedly make use of the following bounds on $s_{p}$
$\sigma_{\min}^{2}p\leq s_{p}\leq\sum_{i=1}^{p}\operatorname{E}\left[\,Y_{i}^{2}\,\right]+\operatorname{E}\left[\,\sum_{i=1}^{p}2|Y_{i}\cdot v_{p,i}|\,\right]+\|\mathbf{v}_{p}\|^{2}\leq\\ Cp+\operatorname{E}\left[\,\sum_{i=1}^{p}|Y_{i}|^{2}+(v_{p,i})^{2}\,\right]+C_{2}p\leq 2(C+C_{2})p,$ (7)
where the lower bound follows from $\operatorname{E}\left[\,(Y_{i}-v_{p,i})^{2}\,\right]\geq\operatorname{E}\left[\,(Y_{i}-\operatorname{E}\left[\,Y_{i}\,\right])^{2}\,\right]=\sigma_{\min}^{2}$, and we used Cauchy’s inequality $|ab|\leq\tfrac{a^{2}}{2}+\tfrac{b^{2}}{2}$.
For the first step of the proof, recall that the expectation of a nonnegative random variable $Y$ can be written as $\int_{0}^{\infty}\operatorname{P}\left(\,Y\geq t\,\right)\mathrm{d}t$, irrespective of whether it is finite or not. Thus
$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]=\int_{0}^{\infty}\operatorname{P}\left(\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\geq t\,\right)\mathrm{d}t=\int_{0}^{\infty}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t.$ (8)
We split the integral in (8) into three parts, obtained by splitting the positive reals at the positions $r^{k}\leq R^{k}$, where
$r$ $:=48(C+C_{2})/\sigma_{\min}^{2}$ $R$ $:=r\cdot p^{6}.$ (9)
We bound the integrand in three different ways, depending on the segment.
Step 1: upper most segment Let $t\in(R^{k},\infty)$, meaning we need to bound the probability that $S(\mathbf{v}_{p})$ is close to zero. Since zero has a finite distance to $\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$, common moment-based concentration inequalities result in a constant bound as $t\to\infty$. This is prohibits their application here, since we need an integrable bound. Instead, we use the existence of densities and their uniform boundedness to obtain a bound that vanishes as $t\to\infty$.
Note that if $X$ has density $f$, then, for arbitrary $m\in\mathbb{R}$, the density of $(X-m)^{2}$ is $0.5(f(m+\sqrt{x})+f(m-\sqrt{x}))/\sqrt{x}$ with $x\in\mathbb{R}_{+}$. Consequently, for all $i\in\mathbb{N},m\in\mathbb{R}$ and $\tau\geq 0$,
$\operatorname{P}\left(\,(Y_{i}-m)^{2}\leq\tau\,\right)\leq\int_{0}^{\tau}\frac{c}{\sqrt{x}}\mathrm{d}x=2c\sqrt{\tau}.$ (10)
We aggregate this bound on an independent subset of $(Y_{i})_{i=1,\ldots,p}$. Note that by construction of $n$, it holds that $(n-1)(M+1)+1\leq p$. By $M$-dependence, $(Y_{j(M+1)+1})_{j=0,\ldots,n-1}$ are independent random variables. Thus
[…]
Indeed, this bound is integrable in $t$:
$\int_{R^{k}}^{\infty}t^{-\frac{n}{2k}}\mathrm{d}t=\frac{1}{-\frac{n}{2k}+1}\left[t^{-\frac{n}{2k}+1}\right]_{R^{k}}^{\infty}\overset{\frac{-n}{2k}+1<0}{=}\\ \frac{1}{\frac{n}{2k}-1}R^{-\frac{n}{2}+k}=\frac{1}{\frac{n}{2k}-1}p^{-3n+6k}r^{-\frac{n}{2}+k}\overset{n\geq 4k}{\leq}p^{-3n+6k}r^{-\frac{n}{2}+k}.$
Thus, the contribution of the upper most segment to the expectation of interest is
[…]
Step 2: middle segment Let $t\in(r^{k},R^{k})$. Then,
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\leq\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right).$ (13)
The right hand side probability, i.e. that a sum of random variables deviates from its expectation, can efficiently be treated by concentration inequalities from the literature. Here, we apply a Hoeffding-type bound allowing for partial dependence among the summands. Since Hoeffding requires bounded random variables, we begin by approximating $S(\mathbf{v}_{p})$ and $s_{p}$ by quantities involving bounded random variables.
According to Lemma A.3, for $B:=\max(\sqrt{2C_{2}},\bar{B})$ with some $\bar{B}$ only depending on global constants, it holds that
$\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\frac{\sigma_{\min}^{2}}{2}$
for all $1\leq i\leq p$. We set $\mathcal{I}:=\{1\leq i\leq p\mid|v_{p,i}|\leq B\}$. Note that $\mathcal{I}$ contains at least $\lfloor\tfrac{p}{2}\rfloor$ elements. Otherwise,
$\|\mathbf{v}_{p}\|^{2}\geq\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}v_{p,i}^{2}>\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}(\sqrt{2C_{2}})^{2}\geq\lceil\frac{p}{2}\rceil 2C_{2}\geq C_{2}p,$
a contradiction to the theorem assumptions. Defining
$A_{p,i}$ $:=(\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)-v_{p,i})^{2}$ $a_{p}$ $:=\operatorname{E}\left[\,\sum_{i\in\mathcal{I}}A_{p,i}\,\right],$
we note that by the previous discussion
$a_{p}\geq\sum_{i\in\mathcal{I}}\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\lfloor\tfrac{p}{2}\rfloor\cdot\frac{\sigma_{\min}^{2}}{2}.$ (14)
Note that for $i\in\mathcal{I}$, we have $A_{p,i}\leq(Y_{i}-v_{p,i})^{2}$. Consequently, $S(\mathbf{v}_{p})\geq\sum_{i\in\mathcal{I}}(Y_{i}-v_{p,i})^{2}\geq\sum_{i\in\mathcal{I}}A_{p,i}$ and we have
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq\frac{s_{p}}{r}\,\right)=\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-(a_{p}-\frac{s_{p}}{r})\,\right).$ (15)
We have
[…]
Recalling $r=48(C+C_{2})/\sigma_{\min}^{2}$, we have
[…]
If $p=2$, then also $\tfrac{2(C+C_{2})p}{r}\leq\lfloor\tfrac{p}{2}\rfloor\tfrac{1}{2}\tfrac{\sigma_{\min}^{2}}{2}\leq\tfrac{a_{p}}{2}$. Thus, in particular for the $p$ fixed at the beginning of the proof, by (16) it holds that
$a_{p}-\frac{s_{p}}{r}\geq\frac{a_{p}}{2}\geq\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}.$
Inserting this in (15), we obtain
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right).$ (17)
We now apply the Hoeffding-type bound. First, note that for $i\in\mathcal{I}$, we have $0\leq A_{p,i}\leq(2B)^{2}$ because both $|\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)|\leq B$ and $|v_{p,i}|\leq B$.
The Hoeffding-type bound from [Janson2004PartlyDependentHoeffding, Thm. 2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$. By $M$-dependence, we may cover $\{1,\ldots,p\}$ by $M+1$ sets of independent variables: consider the index sets $\mathcal{I}_{j}:=\{j+l(M+1):l=0,\ldots,\big{\lfloor}\frac{p-j}{M+1}\big{\rfloor}\}.$ for $j=1,\ldots,M+1$. Then $(A_{p,i})_{i\in\mathcal{I}_{j}}$ are independent for all fixed $j=1,\ldots,M+1$. Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [Janson2004PartlyDependentHoeffding]. Thus
$\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right)\leq\exp\left(-2\frac{(\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4})^{2}}{(M+1)\sum_{i\in\mathcal{I}}(2B)^{4}}\right)\leq\\ \exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right).$
Inserting this into equation (17), we can bound the contribution of the middle segment to the expectation of interest as
[…]
Step 3: lower segment Let $t\in(0,r^{k})$. Since $r^{k}$ is finite and does not depend on $p$, it suffices to bound the integrated probability by $1$. Thus, the contribution of the lower segment to the expectation of interest is
$\int_{0}^{r^{k}}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t\leq r^{k}.$ (19)
Final step It remains to aggregate the three segments. Inserting (12), (18), (19) into (8), we have shown
$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]\leq r^{k}\left(\frac{2c\sqrt{2(C+C_{2})}}{p\sqrt{r}}\right)^{n}p^{-3n/2+6k}+\\ r^{k}\cdot\exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right)\cdot p^{6k}+r^{k}.$
We have shown $\sup_{p}\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]<\infty$ if the right hand side remains bounded as $p\to\infty$. Indeed, the basis of the first term is smaller than $1$ if $p$ is large enough and the exponent is $n\geq 4k>0$; further by $n\geq 4k$ we have $3n/2+6k\leq 0$, and finally the exponential decay by $\lfloor\tfrac{p}{2}\rfloor^{2}/p$ is faster than the growth of $p^{6k}$.
Note that the right hand side of the bound directly only depends on $c,C$ and $C_{2}$, $k$, $\sigma_{\min}^{2}$ and $M$ and that indirectly $r,n,B$ are functions of $M,q,p,C,C_{2},\sigma_{\min}^{2}$ only. Thus, the finite supremum of the right hand side also only depends on those quantities. ∎
Lemma A.3.
Let $(Y_{i})_{i\in\mathbb{N}}$ be a sequence of real random variables such that $\sup_{i\in\mathbb{N}}\operatorname{E}\left[\,|Y_{i}|^{q}\,\right]\leq C_{1}<\infty$ for some $q>2$ and $\inf_{i\in\mathbb{N}}\operatorname{Var}\left[\,Y_{i}\,\right]\geq\sigma_{\min}^{2}$. Then there exists $\bar{B}>0$ depending only $q,C_{1}$ and $\sigma_{\min}^{2}$ such for all $B\geq\bar{B}$
$\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]\geq\frac{\sigma_{\min}^{2}}{2}$
for all $p\in\mathbb{N}$ and $1\leq i\leq p$.
Proof.
First, recall that for any nonnegative random variable $Y$ and $j>0$, we have $\operatorname{E}\left[\,Y^{j}\,\right]=\int_{0}^{\infty}jy^{j-1}\operatorname{P}\left(\,Y>y\,\right)\mathrm{d}y$ [DurrettProbabilityBook, Lemma 2.2.13].
Let $B>0$ be arbitrary. Then, for $j\in\{1,2\}$, $i,p\in\mathbb{N}$, using Markov’s inequality, we have
$\operatorname{E}\left[\,|\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|^{j}\cdot 1_{\{|Y_{i}|>B\}}\,\right]=B^{j}\operatorname{P}\left(\,|Y_{i}|>B\,\right)=\\ B^{j}\operatorname{P}\left(\,|Y_{i}|^{q}>B^{q}\,\right)\leq B^{j}\frac{C_{1}}{B^{q}}=C_{1}B^{j-q}.$
Further, note $\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\cdot 1_{\{|Y_{i}|\leq B\}}=Y_{i}\cdot 1_{\{|Y_{i}|\leq B\}}$. Thus, we have
$\left|\operatorname{Var}\left[\,Y_{i}\,\right]-\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]\right|=|\operatorname{E}\left[\,Y_{i}^{2}\,\right]-\operatorname{E}\left[\,Y_{i}\,\right]^{2}-\\ \operatorname{E}\left[\,(\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B))^{2}\,\right]+\operatorname{E}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]^{2}|\leq\\ \operatorname{E}\left[\,(\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B))^{2}\cdot 1_{\{|Y_{i}|>B\}}\,\right]+\\ \operatorname{E}\left[\,|\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|\cdot 1_{\{|Y_{i}|>B\}}\,\right]\cdot\operatorname{E}\left[\,|Y_{i}+\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|\,\right]\leq\\ C_{1}B^{2-q}+C_{1}B^{1-q}\cdot(C_{1}^{1/q}+B)=2C_{1}B^{2-q}+C_{1}^{1+1/q}B^{1-q}.$
Letting $B\to\infty$, we obtain that the difference of the variances is less than $\tfrac{\sigma_{\min}^{2}}{2}$ for $B$ large enough, which, together with the fact that the bound in the preceeding display is monotonically decreasing in $B$, yields the claim. ∎
See 2.4
Proof.
First, consider $p\geq 4(M+1)+1$ and $1\leq i\leq p$. Note that by Theorem 2.3, $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{-2}\,\right]$ exists, which makes the expectations in this proof well-defined.
Taking (1) componentwise, we find
$|\mu_{i}-m_{p,i}|\leq\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left|\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]\right|.$ (20)
Applying Jensen’s inequality with the convex function $1/\sqrt{x}$, we have
$\frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\leq\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}=1.$
Further, by the Cauchy-Schwarz inequality
[…]
where we applied Theorem 2.3 with $k=1$, using that $\|\mathbf{m}_{p}\|^{2}\leq Cp$ by Lemma 2.2. Inserting both the previous bounds into equation (20) yields the claim for $p\geq 4(M+1)+1$.
When $p<4(M+1)+1$, use Lemma 2.2 and properties of $\mathcal{D}$ to bound
$|\mu_{i}-m_{p,i}|\leq\|\boldsymbol{\mu}-\mathbf{m}_{p}\|\leq\sqrt{Cp}\leq\sqrt{C\cdot(4(M+1)+1)},$
which finalizes the proof. ∎
The following Lemma generalizes Lemma 2.3 to $\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}$ and generalizations thereof. A challenge for this task is that $(\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2})_{p\in\mathbb{N}}$ is not of the form that Lemma 2.3 gives guarantees for, simply because its number of summands does not grow while $|p-i|\leq M$. One can cut out the stagnating region from the sequence (which we we could have done for the asymptotic case $p\gg i$), but then the sequence never contains $\mathbf{m}_{p}$ for $|p-i|\leq M$, which is a problem if one is interested in all $p$. The following Lemma, for some fixed $p$, constructs a sequence admissible for Theorem 2.3 which, for example, agrees with $\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}$ in one index, but may not equal $(\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2})_{p\in\mathbb{N}}$ in other indices.
Definition A.4.
For $i,n\in\mathbb{N}$, define the exclusion sets $\mathcal{E}^{n}(i):=\{i-nM,\ldots,i+nM\}$. We write $\mathcal{E}(i):=\mathcal{E}^{1}(i)$. For $\mathbf{v}\in\mathbb{R}^{p}$, define $\mathbf{v}_{-\mathcal{E}^{n}(i)}:=(v_{j})_{j\in\{1,\ldots,p\}\setminus\mathcal{E}^{n}(i)}$.
Lemma A.5.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, and $k>0,n\in\mathbb{N}$. When $p\geq\lceil 2k\rceil(M+1)+2(nM+1)$, then for any $1\leq i\leq p$
$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{k}}\,\right]\leq\frac{B_{2}(k,n,\mathcal{D})}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}}$ (21)
for a constant $B_{2}<\infty$ only depending on $k,n$ and the $\mathcal{D}$-parameters.
Proof.
Let $p\geq\lceil 2k\rceil(M+1)+2(nM+1)$ and $1\leq i\leq p$ be fixed. Define the exclusion set $\mathcal{E}:=\mathcal{E}^{n}(i)\cap\{1,\ldots,p\}$ and the index set $\check{\mathcal{J}}:=\{1,\ldots,p\}\setminus\mathcal{E}$. By definition, $\mathbf{X}_{-\mathcal{E}^{n}(i)}=(X_{j})_{j\in\check{\mathcal{J}}}$ and $\mathbf{m}_{p,-\mathcal{E}^{n}(i)}=(m_{p,j})_{j\in\check{\mathcal{J}}}$.
Let $\mathcal{J}:=\mathbb{N}\setminus\mathcal{E}$. With $l\in\mathbb{N}$ denoting the sequence index, we plan to apply Theorem 2.3 to $Y_{l}:=X_{\mathcal{J}_{l}}$, where $\mathcal{J}_{l}$ denotes the $l$-th entry of $\mathcal{J}$, and to $\mathbf{v}_{l}:=(m_{l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}\,,\,\mathcal{J}_{j}})_{j=1,\ldots,l}\in\mathbb{R}^{l}$, where $|\cdot|$ denotes set cardinalities and $x_{+}=\max(x,0)$ for $x\in\mathbb{R}$. To see that $\mathbf{v}_{l}$ is well-defined (recall that $m_{a,b}$ requires $b\leq a$), we check $\mathcal{J}_{j}\leq\mathcal{J}_{l}\leq l+|\mathcal{E}|\leq l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}$ for all $1\leq j\leq l$. Regarding the remaining theorem assumptions, note that, just like $(X_{l})_{l}$, we have $(Y_{l})_{l}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$. Further,
[…]
where in the final step we used $1\leq l$.
Thus, applying Theorem 2.3 for $k/2$, with $s_{l}:=\operatorname{E}\left[\,\sum_{j=1}^{l}(Y_{j}-v_{l,j})^{2}\,\right]$, we have
$\sup_{l\geq\lceil 2k\rceil(M+1)+1}\operatorname{E}\left[\,\left(\frac{1}{s_{l}}\sum_{j=1}^{l}(Y_{j}-v_{l,j})^{2}\right)^{-k/2}\,\right]\leq\beta(k/2,2C+C(2nM+1),\mathcal{D})<\infty.$
For translating this result to $\mathbf{X}_{-\mathcal{E}^{n}(i)}$, note that when setting $l:=|\check{\mathcal{J}}|$, we have $(Y_{1},\ldots,Y_{|\check{\mathcal{J}}|})=\mathbf{X}_{-\mathcal{E}^{n}(i)}$ and $\mathbf{v}_{l}=\mathbf{m}_{p,-\mathcal{E}^{n}(i)}$ by $l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}=|\check{\mathcal{J}}|+|\mathcal{E}|+0=p$. Further, using the assumption $p\geq(\lceil 2k\rceil+2n))(M+1)$, we have
$|\check{\mathcal{J}}|=p-|\mathcal{E}|\geq p-(2nM+1)\geq\\ \lceil 2k\rceil(M+1)+2(nM+1)-(2nM+1)=\lceil 2k\rceil(M+1)+1,$
which ensures the sup-constraint is satisfied for $l:=|\check{\mathcal{J}}|$. Thus, in particular we have
$\operatorname{E}\left[\,\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}}{\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{k}}\,\right]\leq\beta(k/2,2C+C(2nM+1),\mathcal{D})<\infty,$
which, after dividing by $\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}\geq(|\check{\mathcal{J}}|\sigma_{\min}^{2})^{k/2}>0$, yields the claim. ∎
Lemma A.6.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$ and $0\leq A,B<\infty$ be constants. Then, for any $i,j\in\mathbb{N}$ we have
$\operatorname{E}\left[\,((|X_{i}|+A)(|X_{j}|+B)^{2})^{q/3}\,\right]\leq B_{4}(A,B,\mathcal{D})<\infty$ (22)
for a constant $B_{4}$ only depending on $A,B$ and the $\mathcal{D}$-parameters
Proof.
First, note that
$(|X_{i}|+A)(|X_{j}|+B)^{2}=|X_{i}|X_{j}^{2}+2|X_{i}||X_{j}|B+|X_{i}|B^{2}+AX_{j}^{2}+2A|X_{j}|B+AB^{2}.$
Noting that by $q\geq 3$ the function $x\mapsto x^{q/3}$ is convex, we have by the convexity definition that
$((|X_{i}|+A)(|X_{j}|+B)^{2})^{q/3}\leq\frac{6^{q/3}}{6}\bigg{(}(|X_{i}|X_{j}^{2})^{q/3}+(2|X_{i}||X_{j}|B)^{q/3}+\\ (|X_{i}|B^{2})^{q/3}+(AX_{j}^{2})^{q/3}+(2A|X_{j}|B)^{q/3}+(AB^{2})^{q/3}\bigg{)}.$ (23)
By Jensen’s inequality, the definition of the $\mathcal{D}$-class (in particular $C\geq 1$), it holds that $\operatorname{E}\left[\,(|X_{i}|^{k})^{q/3}\,\right]\leq\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{k/3}\leq C^{k/3}\leq C$ for all $0<k\leq 2$. The same holds for $|X_{j}|$.
By the Cauchy-Schwarz inequality and Jensen’s inequality, it holds that
$\operatorname{E}\left[\,(|X_{i}||X_{j}|)^{q/3}\,\right]\leq\sqrt{\operatorname{E}\left[\,|X_{i}|^{2q/3}\,\right]}\sqrt{\operatorname{E}\left[\,|X_{j}|^{2q/3}\,\right]}\leq\\ \sqrt{\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{2/3}}\sqrt{\operatorname{E}\left[\,|X_{j}|^{q}\,\right]^{2/3}}\leq C^{1/3}C^{1/3}\leq C.$
Finally, applying Hoelder’s inequality, we find
$\operatorname{E}\left[\,(|X_{i}|X_{j}^{2})^{q/3}\,\right]\leq\||X_{i}|^{q/3}\|_{L^{3}}\||X_{j}|^{2q/3}\|_{L^{3/2}}=\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{1/3}\operatorname{E}\left[\,|X_{j}|^{q}\,\right]^{2/3}\leq C.$
Taking the expectation of (23) and inserting the previous three inequalities yields the claim. ∎
See 2.5
Proof.
First consider $p\geq 14M+8$. Let $1\leq i\leq p$ be fixed and set $\mathcal{E}:=\mathcal{E}^{1}(i)\cap\{1,\ldots,p\}$. Recall $\check{\mathbf{X}}=(X_{i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}}$ and $\check{\mathbf{m}}_{p}=(m_{p,i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}}$. By $M$-dependence of $(X_{i})_{i}$, we have that
$\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]=\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\operatorname{E}\left[\,\frac{1}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]=0,$
since $\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]=0$ and the second factor is finite by Lemma A.5. Thus,
$|m_{p,i}-\mu_{i}|=\left|\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left(\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]-\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]\right)\right|=\\ \left|\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]\right|\leq\\ \operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{|(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}|}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\,\right]=\\ \operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\,\right]\operatorname{E}\left[\,|(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}|\,\right]\leq\\ \frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\cdot\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]^{3/2}}\cdot\\ \operatorname{E}\left[\,\left(\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\right)^{3/2}\,\right]\cdot\sum_{j\in\mathcal{E}}\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|\,\right]\\ =:I\cdot II\cdot III\cdot IV,$
where we used $\sqrt{x}-\sqrt{y}=\tfrac{x-y}{\sqrt{x}+\sqrt{y}}$, followed by the triangle inequality and $\|\mathbf{X}-\mathbf{m}_{p}\|\geq\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\geq\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|\geq 0$, and finally independence of $(X_{j})_{j\in\mathcal{E}}$ and $(X_{j})_{j\in\{1,\ldots,p\}\setminus\mathcal{E}^{2}(i)}$ because of $M$-independence.
For the first Term $I$, it holds that $I\leq 1$ using Jensen’s inequality as shown in the proof of Lemma 2.4.
Regarding Term $II$, note that
$\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}\,\right])+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2},$
since $\operatorname{E}\left[\,(\mathbf{X}-\boldsymbol{\mu})^{T}(\boldsymbol{\mu}-\mathbf{m}_{p})\,\right]=0$, and similarly $\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}_{-\mathcal{E}^{2}(i)}\,\right])+\|\boldsymbol{\mu}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|$. Using Lemma 2.2, properties of $\mathcal{D}$, and $4M+1\leq\tfrac{1}{3}p$ by $p\geq 14M+8$, we have
$II\leq\frac{\sqrt{Cp+Cp}}{((p-(4M+1))\sigma_{\min}^{2}+0)^{3/2}}\leq\frac{\sqrt{p}\sqrt{2C}}{(\tfrac{2}{3}p)^{3/2}\,\sigma_{\min}^{3}}=\frac{1}{p}\cdot\frac{\sqrt{2C}}{\sigma_{\min}^{3}\cdot(\tfrac{2}{3})^{3/2}}.$
Applying Lemma A.5 for $k:=3$ and $n:=2$ to Term $III$, we have $III\leq B_{2}(3,2,\mathcal{D})$.
For Term $IV$, first note $|\mu_{i}|\leq C$ and $|m_{p,j}|\leq B_{1}+C$ according to Lemma 2.4. Thus, applying Jensen’s inequality (note $x\mapsto x^{3/q}$ is concave by $q\geq 3$) and Lemma A.6, we find that for any $j\in\mathcal{E}$ we have
$\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|\,\right]\leq\operatorname{E}\left[\,(|X_{i}|+C)(|X_{j}|+B_{1}+C)^{2}\,\right]\leq\\ \operatorname{E}\left[\,((|X_{i}|+C)(|X_{j}|+B_{1}+C)^{2})^{q/3}\,\right]^{3/q}\leq B_{4}(C,B_{1}+C,\mathcal{D})^{3/q}.$
Since $\mathcal{E}$ contains at most $2M+1$ indices, we have shown that term $IV$ is bounded with constant only depending on the $\mathcal{D}$-parameters.
In summary, Terms $I,III$ and $IV$ are bounded and Term $II$ is $\mathcal{O}(\tfrac{1}{p})$, which yields the Theorem claim for $p\geq 14M+8$.
Let $p\leq 14M+8$. Then, for all $1\leq i\leq p$, by Lemma 2.2,
$|m_{p,i}-\mu_{i}|\leq\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\sqrt{Cp}=\frac{1}{p}\sqrt{C}p^{3/2}\leq\frac{1}{p}\sqrt{C}(14M+8)^{3/2}.$
Taking the maximum of the constants derived in the two cases for $p$ yields the theorem claim. ∎
See 2.6
Proof.
We show
$p\cdot(m_{p,i}-\mu_{i})\xrightarrow{p\to\infty}\frac{1}{2\bar{\sigma}}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right],$
which is equivalent to the claim. Similar to the proof of Theorem 2.5, one can see that for $p\geq i+M$
$p\cdot(m_{p,i}-\mu_{i})=p\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left(\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]-\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]\right)=\\ p\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]=\\ \frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\quad\cdot\quad\frac{p\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{3/2}}\quad\cdot\\ \operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{3/2}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]\\ =:I\cdot II\cdot III.$ (24)
Starting with the first Term, $I$, we note that it can be rewritten as
$I=\operatorname{E}\left[\,\left(\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\right)^{-1/2}\,\right]^{-1}.$
Since by Lemma A.7 we have $\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1$ in probability as $p\to\infty$, also $(\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]})^{-1/2}\to 1$ by the continuos mapping theorem. To show that the random sequence’s expectations also converge, we appeal to a uniform integrability argument [DasGupta2008, Thm 6.2]. Indeed, by Theorem 2.3 it holds for example that
$\sup_{p\geq 4(M+1)+1}\operatorname{E}\left[\,\left(\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\right)^{-1}\,\right]\leq\beta(1,C,\mathcal{D})<\infty.$
[…]
Regarding Term $II$, recall from the proof of Theorem 2.5 that $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}\,\right])+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}$. Further, by Theorem 2.5 it holds that $\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}=\mathcal{O}(\tfrac{1}{p})$ for $p$ large enough. Consequently, $\tfrac{1}{p}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\to\bar{\sigma}^{2}$. Since $\bar{\sigma}\geq\sigma_{\min}>0$, the second Term converges to $\tfrac{1}{\bar{\sigma}^{2}}$.
We treat Term $III$ similar to Term $I$, i.e. first show convergence in probability of the expectation argument and then appeal to a uniform integrability argument to deduce convergence of the expectations. To simplify notation, rewrite $III=\operatorname{E}\left[\,N_{p}/D_{p}\,\right]$, where the numerator $N_{p}$ and denominator $D_{p}$ are given by
$N_{p}$ $:=-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2},$
$D_{p}$ $:=\sqrt{\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\cdot\sqrt{\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\left(\sqrt{\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}+\sqrt{\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\right).$
Starting with the numerator $N_{p}$, note that $m_{p,j}\to\mu_{j}$ by Theorem 2.5. Since $\mathcal{E}(i)$ contains finitely many terms, we conclude by the continuous mapping theorem that $N_{p}\to N:=-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-\mu_{j})^{2}$ in probability (note that for convergence of a random vector in probability it suffices that the components converge in probability).
For studying the denominator $D_{p}$, note that for $p\geq i+M$,
$\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}=\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}-\frac{\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\xrightarrow{p\to\infty}1-0=1$
in probability, since the first term converges to $1$ by Lemma A.7, and the numerator of the second term converges to $\sum_{j\in\mathcal{E}(i)}(X_{j}-\mu_{j})^{2}$ as discussed while $\tfrac{1}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\leq\tfrac{1}{p\sigma_{\min}^{2}}\to 0$, implying the second fraction converges to 0 by Slutsky’s theorem.
Thus, another application of the continuous mapping theorem implies $D_{p}\to 2$ in probability. All in all, $N_{p}/D_{p}\to N/2$ in probability by Slutsky’s theorem, where $N$ is as defined previously.
Next, we show $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$. Crucially $q/3>1$, which is required for the uniform integrability argument. First, note that similar to the proof of Theorem 2.5,
$\operatorname{E}\left[\,\left(\frac{|N_{p}|}{D_{p}}\right)^{q/3}\,\right]\leq\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}\operatorname{E}\left[\,\left(\frac{|N_{p}|}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\right)^{q/3}\,\right]=\\ \frac{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]^{q/2}}\cdot\operatorname{E}\left[\,|N_{p}|^{q/3}\,\right]\cdot\\ \operatorname{E}\left[\,\left(\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\right)^{q/2}\,\right]=:III^{\prime}\cdot III^{\prime\prime}\cdot III^{\prime\prime\prime}.$
We have, using Lemma 2.2, that for $p\geq 4M+2$
$(III^{\prime})^{2/q}=\frac{\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}\,\right]}+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}}{\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}_{-\mathcal{E}^{2}(i)}\,\right]}+\|\boldsymbol{\mu}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\leq\\ \frac{Cp+Cp}{(p-|\mathcal{E}^{2}(i)|)\sigma_{\min}^{2}+0}\leq\frac{2Cp}{(p-(4M+1))\sigma_{\min}^{2}},$
which remains bounded as $p\to\infty$ with bound only depending on the $\mathcal{D}$-parameters.
For the second term, use convexity of $x\mapsto x^{q/3}$, the fact that $|\mu_{i}|\leq C$ and $|m_{p,j}|\leq B_{1}+C$, and Lemma A.6 to find
$III^{\prime\prime}\leq|\mathcal{E}(i)|^{q/3}\sum_{j\in\mathcal{E}(i)}\frac{1}{|\mathcal{E}(i)|}\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|^{q/3}\,\right]\leq\\ |\mathcal{E}(i)|^{(q-3)/3}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,((|X_{1}|+C)(|X_{j}|+B_{1}+C)^{2})^{q/3}\,\right]\leq\\ (2M+1)^{(q-3)/3}(2M+1)B_{4}(C,B_{1}+C,\mathcal{D}).$
Finally, when $p$ large enough, then by Lemma A.5, we have $III^{\prime\prime\prime}\leq B_{2}(q,2,\mathcal{D})$.
In summary, $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$ when $p$ large enough. By [DasGupta2008, Thm 6.2], we conclude $III\to\operatorname{E}\left[\,N/2\,\right]$ as $p\to\infty$.
Inserting the obtained limits for $I$, IIII, and $III$ into (24) yields the claim. ∎
Lemma A.7.
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Then, $\|\mathbf{X}-\mathbf{m}_{p}\|^{2}/\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\to 1$ in probability.
Proof.
If $q\geq 4$, the claim can be shown using Chebychev’s inequality. To allow $q>3$ in Theorem 2.6, we instead employ a truncation approach similar to [DurrettProbabilityBook, Sect. 2.2.3]. For $1\leq i\leq p$, define $A_{p,i}:=(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}$ and set $a_{p}:=\operatorname{E}\left[\,\sum_{i=1}^{p}A_{p,i}\,\right]$.
First, note by the triangle inequality, convexity of $x\mapsto x^{q}$, and Lemma 2.4 that
$\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq 2^{q}(\tfrac{1}{2}\operatorname{E}\left[\,|X_{i}|^{q}\,\right]+\tfrac{1}{2}|m_{p,i}|^{q})\leq 2^{q-1}(C+B_{1}^{q}):=\tilde{C}$ (25)
for $1\leq i\leq p$. Further, for $t>0$ by Markov’s inequality,
$\operatorname{P}\left(\,(X_{i}-m_{p,i})^{2}>t\,\right)=\operatorname{P}\left(\,|X_{i}-m_{p,i}|^{q}>t^{q/2}\,\right)\leq\frac{\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]}{t^{q/2}}\leq\frac{\tilde{C}}{t^{q/2}}.$ (26)
Third, using $q>2$, we make use of the bound
[…]
To prove convergence in probability, let $t>0$. Then,
$\operatorname{P}\left(\,\left|\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{a_{p}}-1\right|>t\,\right)=\operatorname{P}\left(\,\left|\|\mathbf{X}-\mathbf{m}_{p}\|^{2}-a_{p}\right|>a_{p}t\,\right)\leq\\ \operatorname{P}\left(\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\neq\sum_{i=1}^{p}A_{p,i}\,\right)+\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right).$ (28)
We have by inclusion of events that by $q>2$,
[…]
We treat the second term with Chebycheff’s inequality. First, when $q\geq 4$, $\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\operatorname{E}\left[\,(X_{i}-m_{p,i})^{4}\,\right]\leq\tilde{C}^{4/q}\leq\tilde{C}$ since $\tilde{C}\geq C\geq 1$ and using Jensen’s inequality. If $q<4$,
$\operatorname{E}\left[\,A_{p,i}^{2}\,\right]=\operatorname{E}\left[\,((X_{i}-m_{p,i})^{2})^{2-q/2}|X_{i}-m_{p,i}|^{q}1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}\,\right]\leq\\ p^{2-q/2}\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq p^{2-q/2}\tilde{C}$.
Thus, in any case, $\operatorname{Var}\left[\,A_{p,i}\,\right]\leq\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\tilde{C}\cdot\max(1,p^{2-q/2})$.
Let $i\in\mathbb{N}$. By $M$-dependence, $\operatorname{Cov}(A_{p,i},A_{p,j})=0$ for all but $2M+1$ indices $j$. This fact together with $|\operatorname{Cov}(X,Y)|\leq\sqrt{\operatorname{Var}\left[\,X\,\right]\operatorname{Var}\left[\,Y\,\right]}$ gives
$\left|\sum_{j=1}^{p}\operatorname{Cov}(A_{p,i},A_{p,j})\right|\leq(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2}).$
Thus,
$\operatorname{E}\left[\,\left(\sum_{i=1}^{p}A_{p,i}-a_{p}\right)^{2}\,\right]=\operatorname{Var}\left[\,\sum_{i=1}^{p}A_{p,i}\,\right]\leq p\cdot(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2})$
Hence, by Chebycheff’s inequality,
$\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right)\leq\frac{p\cdot(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2})}{t^{2}a_{p}^{2}}.$
Note
[…]
Thus, for $p$ large enough,
$\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right)\leq\frac{1}{t^{2}}\cdot\frac{4(2M+1)\tilde{C}}{\sigma_{\min}^{4}}\cdot\max\left(\frac{1}{p},p^{1-q/2}\right)\xrightarrow{p\to\infty}0.$ (30)
Combining (30) with (29) in (28), we have shown
$\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{a_{p}}\xrightarrow{P}1$ (31)
in probability as $p\to\infty$. Further, note
$\frac{\sum_{i=1}^{p}\operatorname{E}\left[\,(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}>p\}}\,\right]}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\leq\frac{p^{2-q/2}}{p\,\sigma_{\min}^{2}}\left(\tilde{C}+\frac{\tilde{C}}{q/2-1}\right)\xrightarrow{p\to\infty}0,$
which implies
$\frac{a_{p}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}=1-\frac{\sum_{i=1}^{p}\operatorname{E}\left[\,(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}>p\}}\,\right]}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\xrightarrow{p\to\infty}1.$
Combining this with (31) yields the claim by Slutsky’s theorem. ∎
A.2 Further analysis of the simulation results
Linear Regression slopes for the simulation curves can be found in Table 1. The Pareto(3.1) results are analyzed in more detail in Figure 2.
Exp[0.75, 1.25] MA(2) Pareto(3.1) Pareto[7,8]
Slope of $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ 0.49-0.49 0.51-0.51 0.28-0.28 0.45-0.45
Slope of Residual 1.45-1.45 1.64-1.64 1.03-1.03 1.67-1.67
Table 1: Linear regression slopes to the simulation curves after $p\geq 100$
Figure 2: A closer look at the Pareto(3.1) simulations. Left: $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ (same as left side of Figure 1) in pink versus $f(p):=\sqrt{p}\cdot|\tfrac{-19.52}{p}+\tfrac{20.16}{p^{1.03}}|$ in black. Indeed, $f$ explains the decay of $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ well. Right: Plot $f(p)$ from 100100 to $10^{5}$ in black and compare with $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ in blue. Even for $p$ so large, $f$ is still not fully parallel to $\mathcal{O}(\tfrac{1}{\sqrt{p}})$.

### v2

@@
 .
-A method to balance robustness and bias is the median-of-means [Lugosi2019, Minsker2015] or “bucketing” approach [Guerraoui2024].
+A method to balance robustness and bias is the median-of-means [8, 10] or “bucketing” approach [5].
 In this paper, we take a step back and quantify the actual bias of the geometric median, focusing on high-dimensional distributions.
@@
 4 Discussion
-The bias of the geometric median was recently highlighted in [Minsker2024GeometricMedian], where a constant upper bound on the Euclidean bias for a large class of distributions is derived.
+The bias of the geometric median was recently highlighted in [11], where a constant upper bound on the Euclidean bias for a large class of distributions is derived.
 We complement this result by showing the Euclidean bias to be $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ for a smaller class.
@@
 From a robustness viewpoint it raises concern, as the mean is not robust against, say, contaminating a portion of the observations.
-A differentiation clarifies this: while $M$-dependence is compatible with certain cellwise contamination where indeed the geometric median is not always robust [RAYMAEKERS2024Cellwise], rowwise contaminated observations are typically not $M$-dependent and the geometric median is robust here [Minsker2015, schwank2025robustscorematching].
-There is empirical evidence that the geometric median may similarly be close to the mean for sparsely interacting random fields [schwank2025robustscorematching].
-Our results link to the classic paper by [Brown1983] showing that (sample) geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
+A differentiation clarifies this: while $M$-dependence is compatible with certain cellwise contamination where indeed the geometric median is not always robust [12], rowwise contaminated observations are typically not $M$-dependent and the geometric median is robust here [10, 13].
+There is empirical evidence that the geometric median may similarly be close to the mean for sparsely interacting random fields [13].
+Our results link to the classic paper by [1] showing that (sample) geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
 Finally, we remark that the expansion in Theorem 2.6 always contains the term $\operatorname{E}\left[\,(X_{i}-\mu_{i})^{3}\,\right]$, i.e., the negative component skewness.
-This aligns with statistical intuition for the univariate median [Hippel2005].
+This aligns with statistical intuition for the univariate median [6].
 Acknowledgements This paper is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research.
@@
 When $p\geq M+2$, then $X_{1}$ and $X_{M+2}$ are independent, which ensures the distribution of $\mathbf{X}$ does not concentrate on a one-dimensional linear subspace.
-By [Milasevic1987], the geometric median is unique.
+By [9], the geometric median is unique.
 ∎
@@
 $|\partial_{i}G(\mathbf{m})|=\left|\frac{(m_{i}-X_{i})}{\|\mathbf{X}-\mathbf{m}\|}\right|\leq\frac{\|\mathbf{X}-\mathbf{m}\|_{1}}{\|\mathbf{X}-\mathbf{m}\|}\leq\sqrt{p}$
-by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [Shorack2017, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
+by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [14, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
 $\nabla\operatorname{E}\left[\,G(\mathbf{m})\,\right]=\operatorname{E}\left[\,\frac{\mathbf{m}-\mathbf{X}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$
@@
 First, note that for $i\in\mathcal{I}$, we have $0\leq A_{p,i}\leq(2B)^{2}$ because both $|\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)|\leq B$ and $|v_{p,i}|\leq B$.
-The Hoeffding-type bound from [Janson2004PartlyDependentHoeffding, Thm.
+The Hoeffding-type bound from [7, Thm.
 2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$.
@@
 Then $(A_{p,i})_{i\in\mathcal{I}_{j}}$ are independent for all fixed $j=1,\ldots,M+1$.
-Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [Janson2004PartlyDependentHoeffding].
+Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [7].
 Thus
@@
 Proof.
-First, recall that for any nonnegative random variable $Y$ and $j>0$, we have $\operatorname{E}\left[\,Y^{j}\,\right]=\int_{0}^{\infty}jy^{j-1}\operatorname{P}\left(\,Y>y\,\right)\mathrm{d}y$ [DurrettProbabilityBook, Lemma 2.2.13].
+First, recall that for any nonnegative random variable $Y$ and $j>0$, we have $\operatorname{E}\left[\,Y^{j}\,\right]=\int_{0}^{\infty}jy^{j-1}\operatorname{P}\left(\,Y>y\,\right)\mathrm{d}y$ [4, Lemma 2.2.13].
 Let $B>0$ be arbitrary. Then, for $j\in\{1,2\}$, $i,p\in\mathbb{N}$, using Markov’s inequality, we have
@@
 Since by Lemma A.7 we have $\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1$ in probability as $p\to\infty$, also $(\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]})^{-1/2}\to 1$ by the continuos mapping theorem.
-To show that the random sequence’s expectations also converge, we appeal to a uniform integrability argument [DasGupta2008, Thm 6.2].
+To show that the random sequence’s expectations also converge, we appeal to a uniform integrability argument [3, Thm 6.2].
 Indeed, by Theorem 2.3 it holds for example that
@@
 Finally, when $p$ large enough, then by Lemma A.5, we have $III^{\prime\prime\prime}\leq B_{2}(q,2,\mathcal{D})$.
-In summary, $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$ when $p$ large enough.
-By [DasGupta2008, Thm 6.2], we conclude $III\to\operatorname{E}\left[\,N/2\,\right]$ as $p\to\infty$.
+In summary, $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$ when $p$ large enough. By [3, Thm 6.2], we conclude $III\to\operatorname{E}\left[\,N/2\,\right]$ as $p\to\infty$.
 Inserting the obtained limits for $I$, IIII, and $III$ into (24) yields the claim. ∎
@@
 If $q\geq 4$, the claim can be shown using Chebychev’s inequality.
-To allow $q>3$ in Theorem 2.6, we instead employ a truncation approach similar to [DurrettProbabilityBook, Sect.
+To allow $q>3$ in Theorem 2.6, we instead employ a truncation approach similar to [4, Sect.
 2.2.3].

### v3

@@
 The geometric median provides a robust alternative to the sample mean, mitigaging the impact of outliers.
-The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which in practice often is the learning target [^1] For special distributions like the Gaussian, geometric median and mean can coincide.
+The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which often is the learning target in practice [^1] For special distributions like the Gaussian, geometric median and mean can coincide.
 .
@@
 We prove, when each distribution component only depends on a fixed number of neighboring components, that the maximum-norm distance between geometric median and mean vector is of order $\mathcal{O}(\tfrac{1}{p})$, where $p$ denotes the dimension.
-Furthermore, we show that this rate is sharp in general.
-The results are discussed at the end of the paper. Proofs are deferred to the appendix.
+Furthermore, we derive an explicit expansion of the geometric median components when $p$ tends to infinity, in particular showing $\mathcal{O}(\tfrac{1}{p})$ is sharp in general.
+We discuss connections to the literature and implications of our results at the end of the paper. Proofs are deferred to the appendix.
+2 Background and notation
+Let $\mathbf{X}:=(X_{1},\ldots,X_{p})\in\mathbb{R}^{p}$ be a $p$-variate random vector, and $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ be independent samples from $\mathbf{X}$.
+We denote the Euclidean norm by $\|\cdot\|$.
+Consider the two objective functions
+$D(\mathbf{m}):=\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ and $\hat{D}(\mathbf{m}):=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{x}_{i}-\mathbf{m}\|.$
+Any vector $\mathbf{m}\in\mathbb{R}^{p}$ minimizing $D$ or $\hat{D}$ is called population or sample geometric median of $\mathbf{X}$, respectively.
+A minimizer exists under mild conditions; in fact, by replacing $D(\mathbf{m})$ with the equivalent objective $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|-\|\mathbf{X}\|\,\right]$, moment conditions on $\mathbf{X}$ can be avoided.
+When the minimizer is unique, we write $\mathbf{m}$ for the population and $\hat{\mathbf{m}}$ for the sample geometric median.
+By translation invariance of $D$, one can shift $\mathbf{m}$ to the origin without loss of generality, however we keep $\mathbf{m}$ as a variable to make the distance between median and mean recognizable in the formulas.
+The geometric median $\hat{\mathbf{m}}$ has light tails even under minimal moment assumptions on $\mathbf{X}$ [12].
+This property makes it popular in heavy-tailed location testing where it can help outperform alternatives like Hotelling’s $T^{2}$ [6, 5], and more generally in robust statistics and machine learning [7].
+We write $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ for the population mean and $\operatorname{sd}\left[\,X_{i}\,\right]$ and $\operatorname{Var}\left[\,X_{i}\,\right]$ for the standard deviation and variance of component $X_{i}$.
+For $\mathbf{v}\in\mathbb{R}^{p}$, we let $\|\mathbf{v}\|_{\infty}:=\max(|v_{1}|,\ldots,|v_{p}|)$.
+We use the Landau symbols $\mathcal{O}(\cdot)$ and $o(\cdot)$.
 3 Theoretical results
-We consider random vectors $\mathbf{X}:=(X_{1},\ldots,X_{p})\in\mathbb{R}^{p}$ for $p\in\mathbb{N}$, where the sequence of component random variables $(X_{i})_{i\in\mathbb{N}}$ is from the class $\mathcal{D}\equiv\mathcal{D}(M,q,c,\sigma_{\min},C)$ introduced below.
-A geometric median of $\mathbf{X}$ is any vector $\mathbf{m}\in\mathbb{R}^{p}$ minimizing the expectation $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ for the Euclidean norm $\|\cdot\|$.
-By properties of the class $\mathcal{D}$, there exists a unique geometric median (Lemma A.1), which we denote by $\mathbf{m}_{p}$.
-We then study $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|$, where $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$.
+This section examines the distance between mean and geometric median and obtains the results highlighted in Table 1. We begin with our distributional assumptions.
+Recall that a sequence of random variables $(X_{i})_{i\in\mathbb{N}}$ is $M$-dependent if $(X_{i})_{i=1}^{k}$ and $(X_{i})_{i=k+M+1}^{\infty}$ are independent for all $k\in\mathbb{N}$.
+$M$-dependence of finite-dimensional random vectors $\mathbf{X}=(X_{i})_{1\leq i\leq p}$ is defined analogously under restriction to subvectors of $\mathbf{X}$.
 Definition 3.1.
-Let $M\geq 0$ be an integer, and let $c,\sigma_{\min}>0$, $q\geq 2$ and $C\geq 1$ be reals.
-The class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ is comprised of all sequences of absolutely continuous r.v.
-$(X_{i})_{i\in\mathbb{N}}$ that are $M$-dependent, with $\operatorname{E}\left[\,|X_{i}|^{q}\,\right]\leq C$, $\operatorname{E}\left[\,|X_{i}-\mu_{i}|^{q}\,\right]\leq C$, and $\operatorname{sd}\left[\,X_{i}\,\right]\geq\sigma_{\min}>0$, and for which the probability density of $X_{i}$ is bounded by $c<\infty$ for all $i\in\mathbb{N}$.
-Recall that $M$-dependence requires independence of $(X_{i})_{i\leq k}$ and $(X_{i})_{i\geq k+M+1}$ for all $k\in\mathbb{N}$.
-Note that $|\mu_{i}|\leq C$ and $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ since $C\geq 1$.
-We build on the following central identity, derived from $\nabla_{\mathbf{m}}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]|_{\mathbf{m}=\mathbf{m}_{p}}=0$ by interchanging derivative with expectation, and then adding $\pm\boldsymbol{\mu}$:
-$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right],$ (1)
-where all expectations exist when $p$ is large enough (Lemma A.2).
-The intuition is now that, by $M$-dependence, $X_{i}-\mu_{i}$ and $\|\mathbf{X}-\mathbf{m}_{p}\|$ are nearly independent for $p$ large, and thus $m_{p,i}-\mu_{i}\approx\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\cdot 1=0$ by (1).
-Formalizing this approach requires controlling $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{-1}\,\right]$, or slight variations involving $\check{\mathbf{X}}$ introduced below, uniformly in the dimension $p$.
-To this end, we combine the two lemmas, where the first is an initial bound.
+Let $M\geq 0$ be an integer, and let $c,\sigma_{\min}>0$, $q\geq 2$, and $C\geq 1$ be reals.
+The distribution of an $M$-dependent random vector $\mathbf{X}=(X_{i})_{1\leq i\leq p}$ or an $M$-dependent sequence $(X_{i})_{i\in\mathbb{N}}$ belongs to the class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ if all components $X_{i}$ admit a Lebesgue density bounded by $c$, with $\operatorname{E}\left[\,|X_{i}|^{q}\,\right]\leq C$ and $\operatorname{sd}\left[\,X_{i}\,\right]\geq\sigma_{\min}>0$.
+Note that the definition implies that $|\mu_{i}|\leq C$ and $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ as $C\geq 1$.
+Result Integrability Reference
+$\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}\leq\text{const}$ $q>2$ Lemma 3.3
+$\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}=\mathcal{O}(\tfrac{1}{p})$ $q\geq 3$ Theorem 3.4
+$m_{p,i}=\mu_{i}-\tfrac{\alpha_{i}}{p}+o(\tfrac{1}{p})$ $q>3$ Theorem 3.5
+Table 1: Theoretical results with required integrability $q$.
+Let the $p$-variate random vector $\mathbf{X}$ belong to the class $\mathcal{D}\equiv\mathcal{D}(M,q,c,\sigma_{\min},C)$.
+Then, $\mathbf{X}$ is guaranteed to have a unique population geometric median $\mathbf{m}$ when $p\geq M+2$ (Lemma A.1).
+We build on the following identity for $\mathbf{m}$, which is obtained by setting derivatives in the geometric median objective to zero:
+$\mathbf{m}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$ (1)
+Equation (1) can be proven rigorously when $p$ is large enough (Lemma A.2).
+The guiding intuition for why $\|\mathbf{m}-\boldsymbol{\mu}\|$ should be small is that, by $M$-dependence, $X_{i}-\mu_{i}$ and $\|\mathbf{X}-\mathbf{m}\|$ are nearly independent for large $p$.
+Were they actually independent, the second expectation in (1) would factorize and imply $\mathbf{m}-\boldsymbol{\mu}=\mathbf{0}$, since $\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]=0$.
+Formalizing this approach requires controlling $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{-1}\,\right]$, or slight variations introduced later, uniformly in the dimension $p$.
+To deal with $\mathbf{m}$ on the right hand side of (1), we use the standard a-priori bound $\|\mathbf{m}-\boldsymbol{\mu}\|^{2}\leq C\,p$ (see Lemma A.3 or Minsker and Strawn 13).
 Lemma 3.2.
-When $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$, then $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\sqrt{Cp}$.
+Let $\mathbf{Y}:=(Y_{1},\ldots,Y_{p})\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, and let $\mathbf{v}\in\mathbb{R}^{p}$ with $\|\mathbf{v}\|^{2}\leq C_{2}\,p$ for some $C_{2}<\infty$.
+When $p\geq\lceil 4k\rceil(M+1)+1$,
+$\operatorname{E}\left[\,\left(\frac{1}{\|\mathbf{Y}-\mathbf{v}\|^{2}}\right)^{k}\,\right]\leq\frac{\beta(k,C_{2},\mathcal{D})}{\operatorname{E}\left[\,\|\mathbf{Y}-\mathbf{v}\|^{2}\,\right]^{k}}<\infty$ (2)
+for any $k>0$ and a constant $\beta$ depending only on $k,C_{2}$ and the $\mathcal{D}$-parameters.
+Applying the Cauchy-Schwarz inequality to the second expectation in (1) and inserting Lemma 3.2 yields the following very useful intermediate bound.
 Lemma 3.3.
-Let $(Y_{i})_{i\in\mathbb{N}}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$ and $\mathbf{v}_{p}\in\mathbb{R}^{p}$ with $\|\mathbf{v}_{p}\|^{2}\leq C_{2}p$ for $p\in\mathbb{N}$ and some $C_{2}<\infty$.
-With $\mathbf{Y}:=(Y_{1},\ldots,Y_{p})$ we have
-$\sup_{p\geq\lceil 4k\rceil(M+1)+1}\operatorname{E}\left[\,\left(\frac{\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}\,\right]}\right)^{-k}\,\right]=\beta(k,C_{2},\mathcal{D})<\infty,$ (2)
-for any $k>0$ and constant $\beta$ depending only on $k,C_{2}$ and the $\mathcal{D}$-parameters.
-Both Lemmas together with the Cauchy-Schwarz inequality applied to the second expectation in (1) lead to the following intermediate bound.
-Lemma 2.4.
-If $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, then for a constant $B_{1}$ only depending on the $\mathcal{D}$-parameters, $\sup_{p\in\mathbb{N};\;1\leq i\leq p}|\mu_{i}-m_{p,i}|\leq B_{1}$.
-We can now derive the main results, which refine the Cauchy-Schwarz bound in the proof of Lemma 2.4.
-More precisely, we show that $\operatorname{E}\left[\,(X_{i}-\mu_{i})/\|\mathbf{X}-\mathbf{m}_{p}\|\,\right]$ is close to zero by inserting $0=\operatorname{E}\left[\,(X_{i}-\mu)/\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\,\right]$, where $\check{\mathbf{X}}$ is missing all components with indices in $i-M,\ldots,i+M$.
-Formally, using the exclusion set $\mathcal{E}(i):=\mathbb{N}\cap\{i-M,\ldots,i+M\}$, we define $\check{\mathbf{v}}:=(v_{i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}(i)}\in\mathbb{R}^{|\{1,\ldots,p\}\setminus\mathcal{E}(i)|}$ for any vector $\mathbf{v}=(v_{i})_{i=1}^{p}\in\mathbb{R}^{p}$.
-Thus, subtracting zero from the right hand side of (1), a bit of algebra yields that for $p\geq i+M$,
-$m_{p,i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right].$ (3)
-The first expectation is $\mathcal{O}(\sqrt{p})$ while the second is $\mathcal{O}(\tfrac{1}{p^{3/2}})$, leading to
+Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$.
+Then, $\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}\leq B_{1}$ for a constant $B_{1}$ depending only on the $\mathcal{D}$-parameters.
+For deriving the main results, we measure the distance between each right hand side component of equation (1) and zero in a special way.
+Precisely, let $i\leq p$ be fixed and define $\check{\mathbf{X}}:=(X_{j})_{\{1\leq j\leq p\,:\,|j-i|>M\}}$.
+With $\check{\mathbf{m}}$ defined similarly, $\operatorname{E}\left[\,\nicefrac{{(X_{i}-\mu_{i})}}{{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}}\,\right]=0$ by $M$-dependence.
+Subtracting this zero from equation (1), we obtain after a bit of algebra that
+$m_{i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}}{\|\mathbf{X}-\mathbf{m}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}\|(\|\mathbf{X}-\mathbf{m}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}\|)}\,\right],$ (3)
+as proven for $p$ large enough in Lemma A.4. The first expectation is $\mathcal{O}(\sqrt{p})$ while the second is $\mathcal{O}(p^{-3/2})$, leading to the first main result:
 Theorem 3.4.
-Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$.
-Then, with $B_{3}$ depending only on the $\mathcal{D}$-parameters, $|m_{p,i}-\mu_{i}|\leq\frac{B_{3}}{p}$ for all $1\leq i\leq p\in\mathbb{N}$.
-In particular, Theorem 2.5 implies $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\tfrac{B_{3}}{\sqrt{p}}$.
-By $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\to 0$, we can identify the limit in probability of the (suitably scaled) expectation arguments in (3).
-A uniform integrability argument (requiring $q>3$) transfers the convergence to the expectations themselves and yields
+Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$.
+Then, $\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}\leq\tfrac{B_{3}}{p}$ for $B_{3}$ depending only on the $\mathcal{D}$-parameters.
+Next, we consider random vectors $\mathbf{X}_{p}\in\mathbb{R}^{p}$ and their medians $\mathbf{m}_{p}$ as $p$ tends to infinity.
+We establish an asymptotic expansion $m_{p,i}=\mu_{i}-\tfrac{\alpha_{i}}{p}+o(\tfrac{1}{p})$ under mild assumptions.
+This shows that the rate obtained in Theorem 3.4 is generally sharp, and through explicit formulas for $\alpha_{i}$ we gain insight into which distributional properties of $\mathbf{X}$ influence $m_{p,i}-\mu_{i}$ most strongly.
+We prove that $p\cdot(m_{p,i}-\mu_{i})$ converges to a constant $\alpha_{i}$ as $p\to\infty$.
+Precisely, the random variables inside the expectations on the right hand side of the scaled equation (3) admit a limit in probability, since $m_{p,i}\to\mu_{i}$ by Theorem 3.4.
+We transfer this convergence in probability to convergence of the expectations via a uniform integrability argument, requiring $q>3$.
 Theorem 3.5.
-Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>3$. Fix $i\in\mathbb{N}$. If $\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\to\bar{\sigma}^{2}\in\mathbb{R}$, then
-$m_{p,i}=\mu_{i}-\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right]+o\left(\frac{1}{p}\right).$ (4)
+Let $(X_{i})_{i\in\mathbb{N}}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>3$ and assume $\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\to\bar{\sigma}^{2}$.
+Fix $i\in\mathbb{N}$.
+Then, $m_{p,i}=\mu_{i}-\tfrac{\alpha_{i}}{p}+o(\tfrac{1}{p})$, where
+$\alpha_{i}:=\frac{1}{2\bar{\sigma}^{2}}\sum_{|j-i|\leq M}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right].$
+The expansion coeffients $\alpha_{i}$ can be computed analytically in many cases.
+When $M=0$, i.e., $(X_{i})_{i}$ are independent, the sum reduces to the skewness of $X_{i}$.
 Note that while $\sigma_{\min}^{2}\leq\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ always has a convergent subsequence, there may be multiple accumulation points.
+In this case, the expansion is satisfied on each convergent subsequence with individual expansion coefficients $\alpha_{i}$.
 4 Simulations
-We simulate $\hat{\mathbf{m}}_{p}$ for a range of distributions to validate Theorems 2.5 and 2.6. The simulation details are discussed at the end of this section.
+We simulate $\hat{\mathbf{m}}_{p}$ for a range of distributions to validate Theorems 3.4 and 3.5. The simulation details are discussed at the end of this section.
 Figure 1: (Left) Euclidean distance $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$.
@@
 We plot $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$, for which Theorem 3.4 predicts $\mathcal{O}(\tfrac{1}{\sqrt{p}})$, on the left of Figure 1. On the right, we plot the residual
-$\left|\hat{m}_{p,1}-\mu_{1}+\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(1)}\operatorname{E}\left[\,(X_{1}-\mu_{1})(X_{j}-\mu_{j})^{2}\,\right]\right|,$ (5)
+$\big|\hat{m}_{p,1}-\mu_{1}+\frac{1}{2\bar{\sigma}^{2}p}\sum_{j=1}^{M}\operatorname{E}\left[\,(X_{1}-\mu_{1})(X_{j}-\mu_{j})^{2}\,\right]\big|,$ (4)
 for which Theorem 3.5 predicts $o(\tfrac{1}{p})$.
@@
 Table 2 lists linear regression slopes to the simulation curves after $p\geq 100$.
-The simulation study clearly supports the theory, although $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ approaches $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ rather slowly for the heavy-tailed Pareto distributions.
-To see why, note $m_{p,1}-\mu_{1}\approx\tfrac{-19.5}{p}+\tfrac{20.2}{p^{1.03}}$ for Pareto(3.1) from the component expansion plot.
+The simulation study clearly supports the theoretical predictions.
+For the heavy tailed Pareto distributions, $\|\hat{\mathbf{m}}_{p}-\boldsymbol{\mu}\|$ approaches $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ rather slowly in the left hand plot.
+To see why, note $m_{p,1}-\mu_{1}\approx\tfrac{-19.5}{p}+\tfrac{20.2}{p^{1.03}}$ for Pareto(3.1) from the component expansion plot on the right.
 Since all geometric median components are equal in the iid setting, $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\approx\sqrt{p}\cdot|\tfrac{-19.5}{p}+\tfrac{20.2}{p^{1.03}}|$, which indeed fits the simulations results (Figure 2).
-The near cancellation in the absolute value slows the decay for moderate $p$.
-[…]
+The near cancellation considerably slows the decay for moderate dimension $p$.
+To ensure that $\hat{\mathbf{m}}_{p}$ is sufficiently close to $\mathbf{m}_{p}$, we compute $\hat{\mathbf{m}}_{p}$ from $n:=p^{3}$ samples.
+Then, for a large class of distributions and $s>0$, we have $\|\hat{\mathbf{m}}_{p}-\mathbf{m}_{p}\|\lesssim\sqrt{\tfrac{p}{n}}=\tfrac{1}{p}$ with probability $1-e^{4s}$ up to constants not depending on $p,n$ [13, Thm.
+3.9].
+While $n:=p^{3}$ controls the simulation error, it introduces a memory bottleneck.
+Therefore, we compute $\hat{\mathbf{m}}_{p}$ by the online averaged Robbins-Monro algorithm with the default parameter choices $c_{\gamma}=2$ and $\alpha=\tfrac{3}{4}$ [2].
 5 Discussion
-The bias of the geometric median was recently highlighted in [11], where a constant upper bound on the Euclidean bias for a large class of distributions is derived.
+The bias of the geometric median was recently highlighted by Minsker and Strawn [13], where a constant upper bound on the Euclidean bias for a large class of distributions is derived.
 We complement this result by showing the Euclidean bias to be $\mathcal{O}(\tfrac{1}{\sqrt{p}})$ for a smaller class.
-From an estimation viewpoint, a small distance to the mean is desirable.
-From a robustness viewpoint it raises concern, as the mean is not robust against, say, contaminating a portion of the observations.
-A differentiation clarifies this: while $M$-dependence is compatible with certain cellwise contamination where indeed the geometric median is not always robust [12], rowwise contaminated observations are typically not $M$-dependent and the geometric median is robust here [10, 13].
-There is empirical evidence that the geometric median may similarly be close to the mean for sparsely interacting random fields [13].
-Our results link to the classic paper by [1] showing that (sample) geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
-Finally, we remark that the expansion in Theorem 2.6 always contains the term $\operatorname{E}\left[\,(X_{i}-\mu_{i})^{3}\,\right]$, i.e., the negative component skewness.
+Our expansion in Theorem 3.5 always contains the term $\operatorname{E}\left[\,(X_{i}-\mu_{i})^{3}\,\right]$, i.e., the negative component skewness.
 This aligns with statistical intuition for the univariate median [8].
+Our result also links to the classic paper of [1] showing that sample geometric median and mean are equivalent in terms of asymptotic variance for isotropic Gaussians as the ambient dimension approaches infinity.
+Theorem 3.5 can also be applied to the geometric median of means, i.e., when $X_{i}=\tfrac{1}{K}\sum_{k=1}^{K}X_{i}^{(k)}$ with the sequences $(X_{i})_{i\in\mathbb{N}}^{(k)}$ for $k=2,\ldots,K$ being iid copies of $(X_{i})_{i\in\mathbb{N}}^{(1)}$.
+In this case, $\alpha_{i}=\nicefrac{{\alpha^{(1)}_{i}}}{{K}}$.
+In other words, the first order bias of the population geometric median of means scales exactly as $\tfrac{1}{K}$ with the block size $K$.
+Understanding the population bias is one step towards characterizing the bias of the sample geometric median of means.
+The geometric median exhibits a benign bias for high-dimensional distributions beyond those considered in this paper, as empirically observed in Schwank et al.
+[15].
+Towards extending the approach of this paper, we note that if through appropriate mixing conditions on $(X_{i})_{i\in\mathbb{N}}$ one can show the convergence in probability $\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1$ similar to Lemma A.7, the right hand side components of equation (1) converge to $1\cdot\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]=0$, provided one can exchange limit and expectation under a uniform integrability argument similar to the proof of Theorem 3.5.
+This would indeed imply that $m_{p,i}$ converges to $\mu_{i}$ in more general settings.
+Finally, we would like to emphasize that our theory supports existing literature by highlighting that robustness of the geometric median may hinge on the dependence structure of contamination patterns.
+Consider $X_{i}=(1-d_{i})\cdot X_{0,i}+d_{i}\cdot Z_{i}$, where $\mathbf{d}\in\{0,1\}^{p}$ is a random vector independent of the rest.
+The mean is said to not be robust against this contamination, since $\mu_{i}=\operatorname{P}\left(\,d_{i}=0\,\right)\mu_{0,i}+\operatorname{P}\left(\,d_{i}=1\,\right)\operatorname{E}\left[\,Z_{i}\,\right]$ depends strongly on $\operatorname{E}\left[\,Z_{i}\,\right]$.
+When $\mathbf{d}$ is $M$-dependent, our theory applies under appropriate conditions on $\mathbf{X}_{0}$ and $\mathbf{Z}$, suggesting the geometric median need not be robust in such a setting.
+This is consistent with the observations of Raymaekers and Rousseeuw [14], as $M$-dependence of $\mathbf{d}$ is a form of cellwise contamination.
+On the other hand, when $\mathbf{d}$ is fully dependent, i.e., all $d_{i}$ are almost surely equal, also called rowwise contamination, the geometric median is provably robust [12].
+Indeed, our theory does not apply in this case due to lacking $M$-dependence.
 Acknowledgements This paper is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research.
@@
 The authors of this work take full responsibility for its content.
-Appendix A Appendix
-A.1 Proofs
+Supplementary Material
+Appendix A Proofs and additional Lemmas
+A.1 Proof of Theorem 3.4
+See 3.4
+Proof.
+First, consider the case $p\geq 10M+8$. Let $i\in\{1,\ldots,p\}$. Note that
+$\operatorname{E}\big[|(X_{i}-\mu_{i})\sum_{|i-j|\leq M}(X_{j}-m_{j})^{2}|\big]\leq B$ (5)
+for a constant $B$ only depending on the $\mathcal{D}$-parameters (see Lemma A.6).
+We consider indices with sufficient distance to $i$, namely $\mathcal{I}:=\{j\in\{1,\ldots,p\}:|j-i|>2M\}$.
+For $\mathbf{v}\in\mathbb{R}^{p}$, let $\mathbf{v}^{*}:=(v_{j})_{j\in\mathcal{I}}\in\mathbb{R}^{d}$, where $d:=|\mathcal{I}|$.
+Lemma A.4 implies that
+$|m_{i}-\mu_{i}|\leq B\cdot\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{1}{2\|\mathbf{X}^{*}-\mathbf{m}^{*}\|^{3}}\,\right],$ (6)
+where we bounded $\|\mathbf{X}-\mathbf{m}\|\geq\|\check{\mathbf{X}}-\check{\mathbf{m}}\|\geq\|\mathbf{X}^{*}-\mathbf{m}^{*}\|$ and used that $\mathbf{X}^{*}$ is independent of $(X_{j})_{|i-j|\leq M}$.
+Applying Jensen’s inequality with the convex function $1/\sqrt{x}$, we have that
+$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\leq\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]}.$ (7)
+Next, apply Lemma 3.2 to $\mathbf{Y}:=\mathbf{X}^{*}$ and $\mathbf{v}:=\mathbf{m}^{*}$ with $k:=\tfrac{3}{2}$.
+This is possible since $\|\mathbf{m}^{*}\|^{2}\leq\sum_{j\in\mathcal{I}}(|\mu_{j}|+|\mu_{j}-m_{j}|)^{2}\leq(C+B_{1})^{2}\cdot d$ by Lemma 3.3, and since $d\geq p-(4M+1)\geq 6(M+1)+1$ by assumption on $p$.
+The resulting bound is:
+$\operatorname{E}\left[\,\frac{\operatorname{E}\left[\,\|\mathbf{X}^{*}-\mathbf{m}^{*}\|^{2}\,\right]^{3/2}}{\|\mathbf{X}^{*}-\mathbf{m}^{*}\|^{3}}\,\right]\leq\beta(\tfrac{3}{2},(B_{1}+C)^{2},\mathcal{D})=:\tilde{B}$.
+(8)
+Combining the bounds and using $\|\boldsymbol{\mu}-\mathbf{m}\|^{2}\leq p\,C$ by Lemma A.3, we find:
+$|m_{i}-\mu_{i}|\leq\frac{B\tilde{B}}{2}\cdot\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}^{*}-\mathbf{m}^{*}\|^{2}\,\right]^{3/2}}\leq\frac{B\tilde{B}}{2}\frac{\sqrt{\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}\,\right]}+\|\boldsymbol{\mu}-\mathbf{m}\|^{2}}}{(\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}^{*}\,\right]}+0)^{3/2}}\\ \leq\frac{B\tilde{B}}{2}\frac{\sqrt{pC+pC}}{((p-(4M+1))\sigma_{\min}^{2})^{3/2}}\leq\frac{1}{p}\cdot\frac{B\tilde{B}\sqrt{2C}}{2\sigma_{\min}^{3}0.5^{3/2}},$
+where we used $4M+1\leq\tfrac{p}{2}$ by the assumptions on $p$. As $i$ was arbitrary, the theorem claim is proven for $p\geq 10M+8$.
+When $p\leq 10M+8$, the claim holds with $B_{3}:=(10M+8)\cdot B_{1}$ by Lemma 3.3. The maximum of both derived constants is a valid choice for $B_{3}$. ∎
+A.2 Proof of Theorem 3.5
+See 3.5
+Proof.
+We show the following identity, which is equivalent to the claim:
+$p\cdot(m_{p,i}-\mu_{i})\xrightarrow{p\to\infty}\frac{-1}{2\bar{\sigma}^{2}}\sum_{|j-i|\leq M}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right].$
+For $p$ large enough, Lemma A.4 implies that
+$p\cdot(m_{p,i}-\mu_{i})=\frac{p\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{3/2}}\,\cdot\,\operatorname{E}\left[\,F_{p}^{-1}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\cdot S_{p}}{F_{p}\cdot\check{F}_{p}\cdot(F_{p}+\check{F}_{p})}\,\right]=:I_{p}\,\cdot\,II_{p},$
+where we defined the two fractions
+$F_{p}$ $:=\sqrt{\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}},$ $\check{F}_{p}$ $:=\sqrt{\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}},$
+and the sum $S_{p}:=\sum_{|j-i|\leq M}(X_{j}-m_{p,j})^{2}$.
+For the first factor $I_{p}$, note $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}\,\right])+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}$.
+Further, $\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}\leq B_{3}^{2}/p$ by Theorem 3.4, implying that
+$I_{p}^{-1}=\frac{1}{p}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\frac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]+\mathcal{O}\left(\frac{1}{p^{2}}\right)\xrightarrow{p\to\infty}\bar{\sigma}^{2}.$
+Consequently, $\lim_{p\to\infty}I_{p}=\tfrac{\bar{\sigma}}{\bar{\sigma}^{3}}=\tfrac{1}{\bar{\sigma}^{2}}$.
+To determine the limit of the second factor $II_{p}$, we adopt a two-step approach.
+In the first step, we determine the limit in probability of the expectation arguments.
+Using Theorem 3.4, the deterministic sequence $(m_{p,j})_{p\in\mathbb{N}}$ converges to $\mu_{j}$ in particular for all $j$ with $|i-j|\leq M$.
+Thus:
+$S_{p}\to\sum_{|j-i|\leq M}(X_{j}-\mu_{j})^{2}=:S$
+in probability as $p\to\infty$. Further, $F_{p}^{2}\to 1$ in probability as shown in Lemma A.7. We conclude $F_{p}=|F_{p}|\to 1$ by the continuous mapping theorem. Finally,
+$\check{F}_{p}^{2}=F_{p}^{2}-S_{p}\cdot\frac{1}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1-S\cdot 0=1$
+by $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{-1}\leq(p\,\sigma_{\min}^{2})^{-1}\to 0$ and Slutsky’s theorem.
+Using the vector continuous mapping theorem and that convergence in probability of a random vector is equivalent to the component convergence, we conclude that
+$\frac{-(X_{i}-\mu_{i})\,S_{p}}{F_{p}\cdot\check{F}_{p}\cdot(F_{p}+\check{F}_{p})}$ $\to\frac{-(X_{i}-\mu_{i})\,S}{2}$ and $F_{p}^{-1}$ $\to 1.$ (9)
+[…]
+$\operatorname{E}\left[\,\left|\frac{(X_{i}-\mu_{i})\,S_{p}}{F_{p}\cdot\check{F}_{p}\cdot(F_{p}+\check{F}_{p})}\right|^{q/3}\,\right]\leq B_{4}(q,\mathcal{D})\cdot\operatorname{E}\left[\,\frac{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}}{(2\|\mathbf{X}^{*}-\mathbf{m}_{p}^{*}\|^{3})^{q/3}}\,\right]\\ \leq\frac{B_{4}\cdot\beta}{2^{q/3}}\frac{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}}{\operatorname{E}\left[\,\|\mathbf{X}^{*}-\mathbf{m}_{p}^{*}\|^{2}\,\right]^{q/2}}\leq\frac{B_{4}\cdot\beta}{2}\left(\frac{2p\,C}{(p-(4M+1))\sigma_{\min}^{2}}\right)^{q/2},$
+where similar to the proof of Theorem 3.4 we used independence of $(X_{i}-\mu_{i})S_{p}$ and $\mathbf{X}^{*}$, then applied Lemma A.6 (introducing $B_{4}$), then Lemma 3.2 applied to $\mathbf{X}^{*}$ (introducing $\beta$), and finally the trace-bound similar to Theorem 3.4.
+The final expression remains bounded as $p\to\infty$, proving uniform integrability with exponent $q/3>1$.
+We may therefore exchange limit and expectation in both terms of $II_{p}$, and the theorem’s claim follows from (9).
+∎
+A.3 Proof of Lemma 3.2
+See 3.2
+Proof.
+Let $k>0$ and $p\geq\lceil 4k\rceil(M+1)+1$.
+To simplify notation and stress the additive structure, we denote $S(\mathbf{v}):=\sum_{i=1}^{p}(Y_{i}-v_{i})^{2}=\|\mathbf{Y}-\mathbf{v}\|^{2}$ and $s:=\operatorname{E}\left[\,S(\mathbf{v})\,\right]$.
+We repeatedly make use of the following bounds on $s$:
+$\sigma_{\min}^{2}p\leq s\leq\sum_{i=1}^{p}2(\operatorname{E}\left[\,Y_{i}^{2}\,\right]+v_{i}^{2})\leq 2(C+C_{2})\,p,$ (10)
+where we used $\operatorname{E}\left[\,(Y_{i}-v_{i})^{2}\,\right]\geq\operatorname{E}\left[\,(Y_{i}-\operatorname{E}\left[\,Y_{i}\,\right])^{2}\,\right]=\sigma_{\min}^{2}$ and Cauchy’s inequality $|ab|\leq\tfrac{a^{2}}{2}+\tfrac{b^{2}}{2}$.
+In particular, it holds that $s>0$ and we therefore may prove the Lemma’s claim by bounding $\operatorname{E}\left[\,(\tfrac{s}{S(\mathbf{v})})^{k}\,\right]$.
+Using $\operatorname{E}\left[\,Z\,\right]=\int_{0}^{\infty}\operatorname{P}\left(\,Z\geq t\,\right)\mathrm{d}t$ for any nonnegative random variable $Z$ (even if the expectation should be infinite), we obtain that
+$\operatorname{E}\left[\,\left(\frac{s}{S(\mathbf{v})}\right)^{k}\,\right]=\int_{0}^{r^{k}}\operatorname{P}\left(\,S(\mathbf{v})\leq s\cdot t^{-1/k}\,\right)\mathrm{d}t+\int_{r^{k}}^{R^{k}}\operatorname{P}\left(\,S(\mathbf{v})\leq s\cdot t^{-1/k}\,\right)\mathrm{d}t\\ +\int_{R^{k}}^{\infty}\operatorname{P}\left(\,S(\mathbf{v})\leq s\cdot t^{-1/k}\,\right)\mathrm{d}t=:I_{1}+I_{2}+I_{3},$ (11)
+where we split the positive reals at $r^{k}\leq R^{k}$ defined by
+$r$ $:=48(C+C_{2})/\sigma_{\min}^{2}$ $R$ $:=r\cdot p^{6}.$ (12)
+The remaining proof bounds the three integrals separately.
+The integral $I_{3}$
+Let $t\in(R^{k},\infty)$, meaning we must bound the probability that $S(\mathbf{v})$ is close to zero. For arbitrary $m\in\mathbb{R}$ and $\tau\geq 0$, we have that
+$\operatorname{P}\left(\,(Y_{i}-m)^{2}\leq\tau\,\right)=\operatorname{P}\left(\,|Y_{i}-m|\leq\sqrt{\tau}\,\right)\leq 2c\sqrt{\tau}.$ (13)
+We aggregate this bound on an independent subset of $(Y_{i})_{i=1,\ldots,p}$ of size $n:=\lfloor\tfrac{p-1}{M+1}\rfloor+1$.
+Note $n\geq 4k$ and $(n-1)(M+1)+1\leq p$ by assumption on $p$.
+By $M$-dependence, $(Y_{j(M+1)+1})_{j=0,\ldots,n-1}$ are independent, and thus
+[…]
+Importantly, $t^{-\tfrac{n}{2k}}$ is integrable on $(R^{k},\infty)$:
+$\int_{R^{k}}^{\infty}t^{-\frac{n}{2k}}\mathrm{d}t=\frac{1}{-\frac{n}{2k}+1}\left[t^{-\frac{n}{2k}+1}\right]_{R^{k}}^{\infty}\overset{\frac{-n}{2k}+1<0}{=}\frac{1}{\frac{n}{2k}-1}R^{-\frac{n}{2}+k}\\ =\frac{1}{\frac{n}{2k}-1}p^{-3n+6k}r^{-\frac{n}{2}+k}\overset{n\geq 4k}{\leq}p^{-3n+6k}r^{-\frac{n}{2}+k}.$
+Thus, we can bound $I_{3}$ as follows:
+[…]
+The integral $I_{2}$
+Let $t\in(r^{k},R^{k})$. Then,
+$\operatorname{P}\left(\,S(\mathbf{v})\leq s\cdot t^{-1/k}\,\right)\leq\operatorname{P}\left(\,S(\mathbf{v})\leq\frac{s}{r}\,\right).$ (16)
+We control the right hand side probability by a Hoeffding-type bound allowing for partial dependence among the summands.
+Since Hoeffding’s inequality requires bounded random variables, we first approximate $S(\mathbf{v})$ and $s$ by quantities involving bounded random variables.
+According to Lemma A.5, for $B:=\max(\sqrt{2C_{2}},\bar{B})$ with some $\bar{B}$ only depending on global constants, it holds that
+$\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\frac{\sigma_{\min}^{2}}{2}$
+for all $1\leq i\leq p$. We set $\mathcal{I}:=\{1\leq i\leq p\mid|v_{i}|\leq B\}$. Note that $\mathcal{I}$ contains at least $\lfloor\tfrac{p}{2}\rfloor$ elements. Otherwise,
+$\|\mathbf{v}\|^{2}\geq\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}v_{i}^{2}>\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}(\sqrt{2C_{2}})^{2}\geq\bigg\lceil\frac{p}{2}\bigg\rceil 2C_{2}\geq C_{2}p,$
+a contradiction to the theorem assumptions. Defining
+$A_{i}$ $:=(\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)-v_{i})^{2},$ $a$ $:=\operatorname{E}\left[\,\sum_{i\in\mathcal{I}}A_{i}\,\right],$
+we conclude by the previous discussion that
+$a\geq\sum_{i\in\mathcal{I}}\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\bigg\lfloor\frac{p}{2}\bigg\rfloor\cdot\frac{\sigma_{\min}^{2}}{2}.$ (17)
+Note that for $i\in\mathcal{I}$, we have $A_{i}\leq(Y_{i}-v_{i})^{2}$. Consequently, $S(\mathbf{v})\geq\sum_{i\in\mathcal{I}}(Y_{i}-v_{i})^{2}\geq\sum_{i\in\mathcal{I}}A_{i}$ and we have that
+$\operatorname{P}\left(\,S(\mathbf{v})\leq\frac{s}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{i}\leq\frac{s}{r}\,\right)=\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{i}\leq a-(a-\frac{s}{r})\,\right).$ (18)
+Using $r=48(C+C_{2})/\sigma_{\min}^{2}$, we find:
+[…]
+Inserting this in (18), we obtain that
+$\operatorname{P}\left(\,S(\mathbf{v})\leq\frac{s}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{i}\leq a-\bigg\lfloor\frac{p}{2}\bigg\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right).$ (19)
+We now apply the Hoeffding-type bound. First, note that for $i\in\mathcal{I}$, we have $0\leq A_{i}\leq(2B)^{2}$ as both $|\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)|\leq B$ and $|v_{i}|\leq B$.
+The Hoeffding-type bound from [9, Thm.
+2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$.
+By $M$-dependence, a valid cover of $\{1,\ldots,p\}$ are the $M+1$ index sets $\mathcal{I}_{j}:=\{j+l(M+1):l=0,\ldots,\big\lfloor\frac{p-j}{M+1}\big\rfloor\}$ for $j=1,\ldots,M+1$.
+Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [9].
+Thus:
+$\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{i}\leq a-\bigg\lfloor\frac{p}{2}\bigg\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right)\leq\exp\left(-2\frac{(\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4})^{2}}{(M+1)\sum_{i\in\mathcal{I}}(2B)^{4}}\right)\\ \leq\exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right).$
+Inserting this into equation (19), we can bound $I_{2}$ as follows:
+[…]
+The integral $I_{1}$
+Bounding the probability term by one, we obtain that
+$I_{1}=\int_{0}^{r^{k}}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t\leq r^{k}.$ (21)
+Final step
+Inserting (15), (20), (21) into (11), we have shown:
+$\operatorname{E}\left[\,\left(\frac{s}{S(\mathbf{v})}\right)^{k}\,\right]\leq r^{k}\left(\frac{2c\sqrt{2(C+C_{2})}}{p\sqrt{r}}\right)^{n}p^{-3n/2+6k}\\ +r^{k}\cdot\exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right)\cdot p^{6k}+r^{k}.$
+Studying the right hand side for large $p$, we note that the fraction in the first term becomes smaller than $1$ and the exponent is $n\geq 4k>0$; further by $n\geq 4k$ we have $3n/2+6k\leq 0$, and finally the exponential decay by $\lfloor\tfrac{p}{2}\rfloor^{2}/p$ is faster than the growth of $p^{6k}$.
+Consequently, the supremum of the right hand side over $p\in\mathbb{N}$ is finite.
+Further, as the right hand side only depends on $c,C,C_{2},k,\sigma_{\min}^{2},M$, and $r,n,B$ are functions of $M,q,p,C,C_{2},\sigma_{\min}^{2}$ only, this finite supremum only depends on those quantities.
+∎
+A.4 Proof of Lemma 3.3
+See 3.3
+Proof.
+First, consider $p\geq 4(M+1)+1$ and $1\leq i\leq p$. Taking the result of Lemma A.2 component-wise, we find:
+$|\mu_{i}-m_{i}|\leq\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\left|\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}\|}\,\right]\right|.$ (22)
+Applying Jensen’s inequality with the convex function $1/\sqrt{x}$, we have that
+$\frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\leq\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]}}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]}}=1.$
+Further, by the Cauchy-Schwarz inequality we find:
+[…]
+where we applied Lemma 3.2 with $k=1$, using that $\|\mathbf{m}\|^{2}\leq 2(C^{2}+C)p$ by Lemma A.3 (see the proof of Lemma A.2).
+Inserting both the previous bounds into equation (22) yields the claim for $p\geq 4(M+1)+1$.
+When $p<4(M+1)+1$, use Lemma A.3 and properties of $\mathcal{D}$ to bound
+$|\mu_{i}-m_{i}|\leq\|\boldsymbol{\mu}-\mathbf{m}\|\leq\sqrt{Cp}\leq\sqrt{C\cdot(4(M+1)+1)},$
+which finalizes the proof. ∎
+A.5 Additional Lemmas
 Lemma A.1.
-Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ and $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$. Then, when $p\geq M+2$, there exists a unique geometric median of $\mathbf{X}$.
+Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$. Then, when $p\geq M+2$, there exists a unique geometric median of $\mathbf{X}$.
 Proof.
@@
 Lemma A.2.
-Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$.
-Set $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$, denote $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ and let $\mathbf{m}_{p}$ be the geometric median of $\mathbf{X}$.
+Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$.
+Let $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ and let $\mathbf{m}$ be the geometric median of $\mathbf{X}$.
 Then, when $p\geq 2(M+1)+1$,
-$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$
+$\mathbf{m}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right]$
 Proof.
-Let $G(\mathbf{m}):=\|\mathbf{X}-\mathbf{m}\|$ for $\mathbf{m}\in\mathbb{R}^{p}$. We have
-$|\partial_{i}G(\mathbf{m})|=\left|\frac{(m_{i}-X_{i})}{\|\mathbf{X}-\mathbf{m}\|}\right|\leq\frac{\|\mathbf{X}-\mathbf{m}\|_{1}}{\|\mathbf{X}-\mathbf{m}\|}\leq\sqrt{p}$
-by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [14, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
+Let $G(\mathbf{m}):=\|\mathbf{X}-\mathbf{m}\|$ for $\mathbf{m}\in\mathbb{R}^{p}$. We have for almost all $\mathbf{m}$ that
+$|\partial_{i}G(\mathbf{m})|=\left|\frac{(m_{i}-X_{i})}{\|\mathbf{X}-\mathbf{m}\|}\right|\leq\frac{\|\mathbf{X}-\mathbf{m}\|}{\|\mathbf{X}-\mathbf{m}\|}=1.$
+Consequently, by [16, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
 $\nabla\operatorname{E}\left[\,G(\mathbf{m})\,\right]=\operatorname{E}\left[\,\frac{\mathbf{m}-\mathbf{X}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$
-Since $\mathbf{m}_{p}$ optimizes $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$, we have
-$0=\operatorname{E}\left[\,\frac{\mathbf{X}-\mathbf{m}_{p}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$ (6)
-By Lemma 2.2 and 2.3 with $k=\tfrac{1}{2}$ (both proven separately later), we have for $p\geq 2(M+1)+1$ that
-$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]\leq\frac{\beta(0.5,C,\mathcal{D})}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\leq\frac{\beta(0.5,C,\mathcal{D})}{\sigma_{\min}\sqrt{p}}<\infty$
-by $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\geq\operatorname{E}\left[\,\|\mathbf{X}-\boldsymbol{\mu}\|^{2}\,\right]$ due to MSE optimality of the mean.
-Therefore, adding $0=\operatorname{E}\left[\,\tfrac{\boldsymbol{\mu}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$ to (6), we find
-$\operatorname{E}\left[\,\frac{\mathbf{m}_{p}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]=\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$
-Pulling constants out of the expectations and dividing by $\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]>0$ (otherwise $\mathbf{X}=\mathbf{m}_{p}$ almost surely, a contradiction to $\sigma_{\min}>0$), we have shown the claim.
+The geometric median $\mathbf{m}$ of $\mathbf{X}$ (see Lemma A.1) optimizes $\mathbf{m}\mapsto\operatorname{E}\left[\,G(\mathbf{m})\,\right]$, implying that
+$0=\operatorname{E}\left[\,\frac{\mathbf{X}-\mathbf{m}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$ (23)
+By Lemma 3.2 with $k=\tfrac{1}{2}$ (note $\|\mathbf{m}\|^{2}\leq 2(\|\boldsymbol{\mu}\|^{2}+\|\mathbf{m}-\boldsymbol{\mu}\|^{2})\leq 2(C^{2}+C)p$ by Lemma A.3; both Lemmas are proven separately later), we have for $p\geq 2(M+1)+1$ that
+$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]\leq\frac{\beta(0.5,2(C^{2}+C),\mathcal{D})}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]}}\leq\frac{\beta(0.5,2(C^{2}+C),\mathcal{D})}{\sigma_{\min}\sqrt{p}}<\infty$
+by $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]\geq\operatorname{E}\left[\,\|\mathbf{X}-\boldsymbol{\mu}\|^{2}\,\right]$ due to MSE optimality of the mean.
+Therefore, adding $0=\operatorname{E}\left[\,\tfrac{\boldsymbol{\mu}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right]$ to (23), we find:
+$\operatorname{E}\left[\,\frac{\mathbf{m}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right]=\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$
+Pulling constants out of the expectations and dividing by $\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]>0$ (otherwise $\mathbf{X}=\mathbf{m}$ almost surely, a contradiction to $\sigma_{\min}>0$), we have shown the claim.
 ∎
-See 2.2
+Lemma A.3.
+Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$. Then, $\|\mathbf{m}-\boldsymbol{\mu}\|\leq\sqrt{Cp}$.
 Proof.
 By assumption, we have $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ for all $i\in\mathbb{N}$.
-Using the triangle inequality, optimality of the geometric median, Jensen’s inequality and the variance bound yields
-$\|\mathbf{m}_{p}-\boldsymbol{\mu}\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\boldsymbol{\mu}\,\right]\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\mathbf{X}\,\right]\|\leq\operatorname{E}\left[\,\|\mathbf{m}_{p}-\mathbf{X}\|\,\right]\leq\\ \operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|\,\right]\leq\sqrt{\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|^{2}\,\right]}\leq\sqrt{Cp}$
+Using the triangle inequality, optimality of the geometric median, Jensen’s inequality and the variance bound yields that
+$\|\mathbf{m}-\boldsymbol{\mu}\|=\|\operatorname{E}\left[\,\mathbf{m}-\boldsymbol{\mu}\,\right]\|=\|\operatorname{E}\left[\,\mathbf{m}-\mathbf{X}\,\right]\|\leq\operatorname{E}\left[\,\|\mathbf{m}-\mathbf{X}\|\,\right]\\ \leq\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|\,\right]\leq\sqrt{\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|^{2}\,\right]}\leq\sqrt{Cp}$
 as desired. ∎
-See 2.3
+Lemma A.4.
+Let the $p$-variate $\mathbf{X}$ belong to the class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$.
+Fix $i\leq p$ and define $\check{\mathbf{X}}:=(X_{j})_{\{1\leq j\leq p\,:\,|j-i|>M\}}$ and similarly $\check{\mathbf{m}}=(m_{j})_{\{1\leq j\leq p\,:\,|j-i|>M\}}$.
+When $p\geq 4(M+1)$, it holds that
+$m_{i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}}{\|\mathbf{X}-\mathbf{m}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}\|(\|\mathbf{X}-\mathbf{m}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}\|)}\,\right].$
 Proof.
-To simplify notation and stress the additive structure, we denote $S(\mathbf{v}_{p}):=\sum_{i=1}^{p}(Y_{i}-v_{p,i})^{2}=\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}$ and $s_{p}:=\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$ for $p\in\mathbb{N}$.
-The quantity of interest thus is $\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]$.
-Let $k>0$ and $p\geq\lceil 4k\rceil(M+1)+1$. We use $n:=\lfloor\tfrac{p-1}{M+1}\rfloor+1$, and note $n\geq 4k$.
-We repeatedly make use of the following bounds on $s_{p}$
-$\sigma_{\min}^{2}p\leq s_{p}\leq\sum_{i=1}^{p}\operatorname{E}\left[\,Y_{i}^{2}\,\right]+\operatorname{E}\left[\,\sum_{i=1}^{p}2|Y_{i}\cdot v_{p,i}|\,\right]+\|\mathbf{v}_{p}\|^{2}\leq\\ Cp+\operatorname{E}\left[\,\sum_{i=1}^{p}|Y_{i}|^{2}+(v_{p,i})^{2}\,\right]+C_{2}p\leq 2(C+C_{2})p,$ (7)
-where the lower bound follows from $\operatorname{E}\left[\,(Y_{i}-v_{p,i})^{2}\,\right]\geq\operatorname{E}\left[\,(Y_{i}-\operatorname{E}\left[\,Y_{i}\,\right])^{2}\,\right]=\sigma_{\min}^{2}$, and we used Cauchy’s inequality $|ab|\leq\tfrac{a^{2}}{2}+\tfrac{b^{2}}{2}$.
-For the first step of the proof, recall that the expectation of a nonnegative random variable $Y$ can be written as $\int_{0}^{\infty}\operatorname{P}\left(\,Y\geq t\,\right)\mathrm{d}t$, irrespective of whether it is finite or not.
-Thus
-$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]=\int_{0}^{\infty}\operatorname{P}\left(\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\geq t\,\right)\mathrm{d}t=\int_{0}^{\infty}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t.$ (8)
-We split the integral in (8) into three parts, obtained by splitting the positive reals at the positions $r^{k}\leq R^{k}$, where
-$r$ $:=48(C+C_{2})/\sigma_{\min}^{2}$ $R$ $:=r\cdot p^{6}.$ (9)
-We bound the integrand in three different ways, depending on the segment.
-Step 1: upper most segment Let $t\in(R^{k},\infty)$, meaning we need to bound the probability that $S(\mathbf{v}_{p})$ is close to zero.
-Since zero has a finite distance to $\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$, common moment-based concentration inequalities result in a constant bound as $t\to\infty$.
-This is prohibits their application here, since we need an integrable bound.
-Instead, we use the existence of densities and their uniform boundedness to obtain a bound that vanishes as $t\to\infty$.
-Note that if $X$ has density $f$, then, for arbitrary $m\in\mathbb{R}$, the density of $(X-m)^{2}$ is $0.5(f(m+\sqrt{x})+f(m-\sqrt{x}))/\sqrt{x}$ with $x\in\mathbb{R}_{+}$.
-Consequently, for all $i\in\mathbb{N},m\in\mathbb{R}$ and $\tau\geq 0$,
-$\operatorname{P}\left(\,(Y_{i}-m)^{2}\leq\tau\,\right)\leq\int_{0}^{\tau}\frac{c}{\sqrt{x}}\mathrm{d}x=2c\sqrt{\tau}.$ (10)
-We aggregate this bound on an independent subset of $(Y_{i})_{i=1,\ldots,p}$.
-Note that by construction of $n$, it holds that $(n-1)(M+1)+1\leq p$.
-By $M$-dependence, $(Y_{j(M+1)+1})_{j=0,\ldots,n-1}$ are independent random variables.
-Thus
-[…]
-Indeed, this bound is integrable in $t$:
-$\int_{R^{k}}^{\infty}t^{-\frac{n}{2k}}\mathrm{d}t=\frac{1}{-\frac{n}{2k}+1}\left[t^{-\frac{n}{2k}+1}\right]_{R^{k}}^{\infty}\overset{\frac{-n}{2k}+1<0}{=}\\ \frac{1}{\frac{n}{2k}-1}R^{-\frac{n}{2}+k}=\frac{1}{\frac{n}{2k}-1}p^{-3n+6k}r^{-\frac{n}{2}+k}\overset{n\geq 4k}{\leq}p^{-3n+6k}r^{-\frac{n}{2}+k}.$
-Thus, the contribution of the upper most segment to the expectation of interest is
-[…]
-Step 2: middle segment Let $t\in(r^{k},R^{k})$. Then,
-$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\leq\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right).$ (13)
-The right hand side probability, i.e.
-that a sum of random variables deviates from its expectation, can efficiently be treated by concentration inequalities from the literature.
-Here, we apply a Hoeffding-type bound allowing for partial dependence among the summands.
-Since Hoeffding requires bounded random variables, we begin by approximating $S(\mathbf{v}_{p})$ and $s_{p}$ by quantities involving bounded random variables.
-According to Lemma A.3, for $B:=\max(\sqrt{2C_{2}},\bar{B})$ with some $\bar{B}$ only depending on global constants, it holds that
-$\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\frac{\sigma_{\min}^{2}}{2}$
-for all $1\leq i\leq p$. We set $\mathcal{I}:=\{1\leq i\leq p\mid|v_{p,i}|\leq B\}$. Note that $\mathcal{I}$ contains at least $\lfloor\tfrac{p}{2}\rfloor$ elements. Otherwise,
-$\|\mathbf{v}_{p}\|^{2}\geq\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}v_{p,i}^{2}>\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}(\sqrt{2C_{2}})^{2}\geq\lceil\frac{p}{2}\rceil 2C_{2}\geq C_{2}p,$
-a contradiction to the theorem assumptions. Defining
-$A_{p,i}$ $:=(\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)-v_{p,i})^{2}$ $a_{p}$ $:=\operatorname{E}\left[\,\sum_{i\in\mathcal{I}}A_{p,i}\,\right],$
-we note that by the previous discussion
-$a_{p}\geq\sum_{i\in\mathcal{I}}\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\lfloor\tfrac{p}{2}\rfloor\cdot\frac{\sigma_{\min}^{2}}{2}.$ (14)
-Note that for $i\in\mathcal{I}$, we have $A_{p,i}\leq(Y_{i}-v_{p,i})^{2}$. Consequently, $S(\mathbf{v}_{p})\geq\sum_{i\in\mathcal{I}}(Y_{i}-v_{p,i})^{2}\geq\sum_{i\in\mathcal{I}}A_{p,i}$ and we have
-$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq\frac{s_{p}}{r}\,\right)=\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-(a_{p}-\frac{s_{p}}{r})\,\right).$ (15)
-We have
-[…]
-Recalling $r=48(C+C_{2})/\sigma_{\min}^{2}$, we have
-[…]
-If $p=2$, then also $\tfrac{2(C+C_{2})p}{r}\leq\lfloor\tfrac{p}{2}\rfloor\tfrac{1}{2}\tfrac{\sigma_{\min}^{2}}{2}\leq\tfrac{a_{p}}{2}$.
-Thus, in particular for the $p$ fixed at the beginning of the proof, by (16) it holds that
-$a_{p}-\frac{s_{p}}{r}\geq\frac{a_{p}}{2}\geq\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}.$
-Inserting this in (15), we obtain
-$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right).$ (17)
-We now apply the Hoeffding-type bound.
-First, note that for $i\in\mathcal{I}$, we have $0\leq A_{p,i}\leq(2B)^{2}$ because both $|\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)|\leq B$ and $|v_{p,i}|\leq B$.
-The Hoeffding-type bound from [7, Thm.
-2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$.
-By $M$-dependence, we may cover $\{1,\ldots,p\}$ by $M+1$ sets of independent variables: consider the index sets $\mathcal{I}_{j}:=\{j+l(M+1):l=0,\ldots,\big\lfloor\frac{p-j}{M+1}\big\rfloor\}.$ for $j=1,\ldots,M+1$.
-Then $(A_{p,i})_{i\in\mathcal{I}_{j}}$ are independent for all fixed $j=1,\ldots,M+1$.
-Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [7].
-Thus
-$\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right)\leq\exp\left(-2\frac{(\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4})^{2}}{(M+1)\sum_{i\in\mathcal{I}}(2B)^{4}}\right)\leq\\ \exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right).$
-Inserting this into equation (17), we can bound the contribution of the middle segment to the expectation of interest as
-[…]
-Step 3: lower segment Let $t\in(0,r^{k})$.
-Since $r^{k}$ is finite and does not depend on $p$, it suffices to bound the integrated probability by $1$.
-Thus, the contribution of the lower segment to the expectation of interest is
-$\int_{0}^{r^{k}}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t\leq r^{k}.$ (19)
-Final step It remains to aggregate the three segments. Inserting (12), (18), (19) into (8), we have shown
-$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]\leq r^{k}\left(\frac{2c\sqrt{2(C+C_{2})}}{p\sqrt{r}}\right)^{n}p^{-3n/2+6k}+\\ r^{k}\cdot\exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right)\cdot p^{6k}+r^{k}.$
-We have shown $\sup_{p}\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]<\infty$ if the right hand side remains bounded as $p\to\infty$.
-Indeed, the basis of the first term is smaller than $1$ if $p$ is large enough and the exponent is $n\geq 4k>0$; further by $n\geq 4k$ we have $3n/2+6k\leq 0$, and finally the exponential decay by $\lfloor\tfrac{p}{2}\rfloor^{2}/p$ is faster than the growth of $p^{6k}$.
-Note that the right hand side of the bound directly only depends on $c,C$ and $C_{2}$, $k$, $\sigma_{\min}^{2}$ and $M$ and that indirectly $r,n,B$ are functions of $M,q,p,C,C_{2},\sigma_{\min}^{2}$ only.
-Thus, the finite supremum of the right hand side also only depends on those quantities.
+Recall that by Lemma A.2 we have:
+$m_{i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$ (24)
+Further, $\check{\mathbf{X}}$ and $\check{\mathbf{m}}$ have at least $2(M+1)+1$ components by assumption on $p$.
+Using Lemma 3.3 to control $\|\check{\mathbf{m}}\|^{2}$, we may apply Lemma 3.2 with $k=\tfrac{1}{2}$ to obtain $\operatorname{E}\left[\,\tfrac{1}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]<\infty$.
+Thus:
+$\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]=\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\operatorname{E}\left[\,\frac{1}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]=0.$
+For reals $x,y>0$ it holds that
+$\frac{1}{\sqrt{x}}-\frac{1}{\sqrt{y}}=\frac{\sqrt{y}-\sqrt{x}}{\sqrt{x}\sqrt{y}}=\frac{y-x}{\sqrt{x}\sqrt{y}\,(\sqrt{x}+\sqrt{y})}.$
+Subtracting $0=\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\tfrac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]$ from the right hand side of (24) and inserting this algebraic relation yields the claim.
 ∎
@@
 Let $(Y_{i})_{i\in\mathbb{N}}$ be a sequence of real random variables such that $\sup_{i\in\mathbb{N}}\operatorname{E}\left[\,|Y_{i}|^{q}\,\right]\leq C_{1}<\infty$ for some $q>2$ and $\inf_{i\in\mathbb{N}}\operatorname{Var}\left[\,Y_{i}\,\right]\geq\sigma_{\min}^{2}$.
-Then there exists $\bar{B}>0$ depending only $q,C_{1}$ and $\sigma_{\min}^{2}$ such for all $B\geq\bar{B}$
+Then there exists $\bar{B}>0$ depending only $q,C_{1}$ and $\sigma_{\min}^{2}$ such that
 $\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]\geq\frac{\sigma_{\min}^{2}}{2}$
-for all $p\in\mathbb{N}$ and $1\leq i\leq p$.
+for all $B\geq\bar{B}$ and $i\in\mathbb{N}$.
 Proof.
 First, recall that for any nonnegative random variable $Y$ and $j>0$, we have $\operatorname{E}\left[\,Y^{j}\,\right]=\int_{0}^{\infty}jy^{j-1}\operatorname{P}\left(\,Y>y\,\right)\mathrm{d}y$ [4, Lemma 2.2.13].
-Let $B>0$ be arbitrary. Then, for $j\in\{1,2\}$, $i,p\in\mathbb{N}$, using Markov’s inequality, we have
+Let $B>0$ be arbitrary. Then, for $j\in\{1,2\}$, $i\in\mathbb{N}$, using Markov’s inequality, we have:
 $\operatorname{E}\left[\,|\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|^{j}\cdot 1_{\{|Y_{i}|>B\}}\,\right]=B^{j}\operatorname{P}\left(\,|Y_{i}|>B\,\right)=\\ B^{j}\operatorname{P}\left(\,|Y_{i}|^{q}>B^{q}\,\right)\leq B^{j}\frac{C_{1}}{B^{q}}=C_{1}B^{j-q}.$
-Further, note $\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\cdot 1_{\{|Y_{i}|\leq B\}}=Y_{i}\cdot 1_{\{|Y_{i}|\leq B\}}$. Thus, we have
+Further, note $\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\cdot 1_{\{|Y_{i}|\leq B\}}=Y_{i}\cdot 1_{\{|Y_{i}|\leq B\}}$. Thus, we have:
 $\left|\operatorname{Var}\left[\,Y_{i}\,\right]-\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]\right|=|\operatorname{E}\left[\,Y_{i}^{2}\,\right]-\operatorname{E}\left[\,Y_{i}\,\right]^{2}-\\ \operatorname{E}\left[\,(\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B))^{2}\,\right]+\operatorname{E}\left[\,\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\,\right]^{2}|\leq\\ \operatorname{E}\left[\,(\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B))^{2}\cdot 1_{\{|Y_{i}|>B\}}\,\right]+\\ \operatorname{E}\left[\,|\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|\cdot 1_{\{|Y_{i}|>B\}}\,\right]\cdot\operatorname{E}\left[\,|Y_{i}+\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)|\,\right]\leq\\ C_{1}B^{2-q}+C_{1}B^{1-q}\cdot(C_{1}^{1/q}+B)=2C_{1}B^{2-q}+C_{1}^{1+1/q}B^{1-q}.$
@@
 ∎
-See 2.4
+Lemma A.6.
+Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$. Then, for all $3\leq k\leq q$, we have that
+$\operatorname{E}\left[\,|\,(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}\,|^{k/3}\,\right]\leq B_{4}(k,\mathcal{D})$
+for all $1\leq i\leq p$ and a constant $B_{4}$ only depending on $k$ and the $\mathcal{D}$-parameters.
 Proof.
-First, consider $p\geq 4(M+1)+1$ and $1\leq i\leq p$.
-Note that by Theorem 2.3, $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{-2}\,\right]$ exists, which makes the expectations in this proof well-defined.
-Taking (1) componentwise, we find
-$|\mu_{i}-m_{p,i}|\leq\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left|\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]\right|.$ (20)
-Applying Jensen’s inequality with the convex function $1/\sqrt{x}$, we have
-$\frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\leq\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}=1.$
-Further, by the Cauchy-Schwarz inequality
-[…]
-where we applied Theorem 2.3 with $k=1$, using that $\|\mathbf{m}_{p}\|^{2}\leq Cp$ by Lemma 2.2. Inserting both the previous bounds into equation (20) yields the claim for $p\geq 4(M+1)+1$.
-When $p<4(M+1)+1$, use Lemma 2.2 and properties of $\mathcal{D}$ to bound
-$|\mu_{i}-m_{p,i}|\leq\|\boldsymbol{\mu}-\mathbf{m}_{p}\|\leq\sqrt{Cp}\leq\sqrt{C\cdot(4(M+1)+1)},$
-which finalizes the proof. ∎
-The following Lemma generalizes Lemma 2.3 to $\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}$ and generalizations thereof.
-A challenge for this task is that $(\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2})_{p\in\mathbb{N}}$ is not of the form that Lemma 2.3 gives guarantees for, simply because its number of summands does not grow while $|p-i|\leq M$.
-One can cut out the stagnating region from the sequence (which we we could have done for the asymptotic case $p\gg i$), but then the sequence never contains $\mathbf{m}_{p}$ for $|p-i|\leq M$, which is a problem if one is interested in all $p$.
-The following Lemma, for some fixed $p$, constructs a sequence admissible for Theorem 2.3 which, for example, agrees with $\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}$ in one index, but may not equal $(\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2})_{p\in\mathbb{N}}$ in other indices.
-Definition A.4.
-For $i,n\in\mathbb{N}$, define the exclusion sets $\mathcal{E}^{n}(i):=\{i-nM,\ldots,i+nM\}$.
-We write $\mathcal{E}(i):=\mathcal{E}^{1}(i)$.
-For $\mathbf{v}\in\mathbb{R}^{p}$, define $\mathbf{v}_{-\mathcal{E}^{n}(i)}:=(v_{j})_{j\in\{1,\ldots,p\}\setminus\mathcal{E}^{n}(i)}$.
-Lemma A.5.
-Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, and $k>0,n\in\mathbb{N}$. When $p\geq\lceil 2k\rceil(M+1)+2(nM+1)$, then for any $1\leq i\leq p$
-$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{k}}\,\right]\leq\frac{B_{2}(k,n,\mathcal{D})}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}}$ (21)
-for a constant $B_{2}<\infty$ only depending on $k,n$ and the $\mathcal{D}$-parameters.
-Proof.
-Let $p\geq\lceil 2k\rceil(M+1)+2(nM+1)$ and $1\leq i\leq p$ be fixed.
-Define the exclusion set $\mathcal{E}:=\mathcal{E}^{n}(i)\cap\{1,\ldots,p\}$ and the index set $\check{\mathcal{J}}:=\{1,\ldots,p\}\setminus\mathcal{E}$.
-By definition, $\mathbf{X}_{-\mathcal{E}^{n}(i)}=(X_{j})_{j\in\check{\mathcal{J}}}$ and $\mathbf{m}_{p,-\mathcal{E}^{n}(i)}=(m_{p,j})_{j\in\check{\mathcal{J}}}$.
-Let $\mathcal{J}:=\mathbb{N}\setminus\mathcal{E}$.
-With $l\in\mathbb{N}$ denoting the sequence index, we plan to apply Theorem 2.3 to $Y_{l}:=X_{\mathcal{J}_{l}}$, where $\mathcal{J}_{l}$ denotes the $l$-th entry of $\mathcal{J}$, and to $\mathbf{v}_{l}:=(m_{l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}\,,\,\mathcal{J}_{j}})_{j=1,\ldots,l}\in\mathbb{R}^{l}$, where $|\cdot|$ denotes set cardinalities and $x_{+}=\max(x,0)$ for $x\in\mathbb{R}$.
-To see that $\mathbf{v}_{l}$ is well-defined (recall that $m_{a,b}$ requires $b\leq a$), we check $\mathcal{J}_{j}\leq\mathcal{J}_{l}\leq l+|\mathcal{E}|\leq l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}$ for all $1\leq j\leq l$.
-Regarding the remaining theorem assumptions, note that, just like $(X_{l})_{l}$, we have $(Y_{l})_{l}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$.
-Further,
-[…]
-where in the final step we used $1\leq l$.
-Thus, applying Theorem 2.3 for $k/2$, with $s_{l}:=\operatorname{E}\left[\,\sum_{j=1}^{l}(Y_{j}-v_{l,j})^{2}\,\right]$, we have
-$\sup_{l\geq\lceil 2k\rceil(M+1)+1}\operatorname{E}\left[\,\left(\frac{1}{s_{l}}\sum_{j=1}^{l}(Y_{j}-v_{l,j})^{2}\right)^{-k/2}\,\right]\leq\beta(k/2,2C+C(2nM+1),\mathcal{D})<\infty.$
-For translating this result to $\mathbf{X}_{-\mathcal{E}^{n}(i)}$, note that when setting $l:=|\check{\mathcal{J}}|$, we have $(Y_{1},\ldots,Y_{|\check{\mathcal{J}}|})=\mathbf{X}_{-\mathcal{E}^{n}(i)}$ and $\mathbf{v}_{l}=\mathbf{m}_{p,-\mathcal{E}^{n}(i)}$ by $l+|\mathcal{E}|+(l-|\check{\mathcal{J}}|)_{+}=|\check{\mathcal{J}}|+|\mathcal{E}|+0=p$.
-Further, using the assumption $p\geq(\lceil 2k\rceil+2n))(M+1)$, we have
-$|\check{\mathcal{J}}|=p-|\mathcal{E}|\geq p-(2nM+1)\geq\\ \lceil 2k\rceil(M+1)+2(nM+1)-(2nM+1)=\lceil 2k\rceil(M+1)+1,$
-which ensures the sup-constraint is satisfied for $l:=|\check{\mathcal{J}}|$. Thus, in particular we have
-$\operatorname{E}\left[\,\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}}{\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{k}}\,\right]\leq\beta(k/2,2C+C(2nM+1),\mathcal{D})<\infty,$
-which, after dividing by $\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{n}(i)}-\mathbf{m}_{p,-\mathcal{E}^{n}(i)}\|^{2}\,\right]^{k/2}\geq(|\check{\mathcal{J}}|\sigma_{\min}^{2})^{k/2}>0$, yields the claim.
-∎
-Lemma A.6.
-Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$ and $0\leq A,B<\infty$ be constants. Then, for any $i,j\in\mathbb{N}$ we have
-$\operatorname{E}\left[\,((|X_{i}|+A)(|X_{j}|+B)^{2})^{q/3}\,\right]\leq B_{4}(A,B,\mathcal{D})<\infty$ (22)
-for a constant $B_{4}$ only depending on $A,B$ and the $\mathcal{D}$-parameters
-Proof.
-First, note that
-$(|X_{i}|+A)(|X_{j}|+B)^{2}=|X_{i}|X_{j}^{2}+2|X_{i}||X_{j}|B+|X_{i}|B^{2}+AX_{j}^{2}+2A|X_{j}|B+AB^{2}.$
-Noting that by $q\geq 3$ the function $x\mapsto x^{q/3}$ is convex, we have by the convexity definition that
-$((|X_{i}|+A)(|X_{j}|+B)^{2})^{q/3}\leq\frac{6^{q/3}}{6}\bigg((|X_{i}|X_{j}^{2})^{q/3}+(2|X_{i}||X_{j}|B)^{q/3}+\\ (|X_{i}|B^{2})^{q/3}+(AX_{j}^{2})^{q/3}+(2A|X_{j}|B)^{q/3}+(AB^{2})^{q/3}\bigg).$ (23)
-By Jensen’s inequality, the definition of the $\mathcal{D}$-class (in particular $C\geq 1$), it holds that $\operatorname{E}\left[\,(|X_{i}|^{k})^{q/3}\,\right]\leq\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{k/3}\leq C^{k/3}\leq C$ for all $0<k\leq 2$.
-The same holds for $|X_{j}|$.
-By the Cauchy-Schwarz inequality and Jensen’s inequality, it holds that
-$\operatorname{E}\left[\,(|X_{i}||X_{j}|)^{q/3}\,\right]\leq\sqrt{\operatorname{E}\left[\,|X_{i}|^{2q/3}\,\right]}\sqrt{\operatorname{E}\left[\,|X_{j}|^{2q/3}\,\right]}\leq\\ \sqrt{\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{2/3}}\sqrt{\operatorname{E}\left[\,|X_{j}|^{q}\,\right]^{2/3}}\leq C^{1/3}C^{1/3}\leq C.$
-Finally, applying Hoelder’s inequality, we find
-$\operatorname{E}\left[\,(|X_{i}|X_{j}^{2})^{q/3}\,\right]\leq\||X_{i}|^{q/3}\|_{L^{3}}\||X_{j}|^{2q/3}\|_{L^{3/2}}=\operatorname{E}\left[\,|X_{i}|^{q}\,\right]^{1/3}\operatorname{E}\left[\,|X_{j}|^{q}\,\right]^{2/3}\leq C.$
-Taking the expectation of (23) and inserting the previous three inequalities yields the claim. ∎
-See 2.5
-Proof.
-First consider $p\geq 14M+8$.
-Let $1\leq i\leq p$ be fixed and set $\mathcal{E}:=\mathcal{E}^{1}(i)\cap\{1,\ldots,p\}$.
-Recall $\check{\mathbf{X}}=(X_{i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}}$ and $\check{\mathbf{m}}_{p}=(m_{p,i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}}$.
-By $M$-dependence of $(X_{i})_{i}$, we have that
-$\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]=\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\operatorname{E}\left[\,\frac{1}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]=0,$
-since $\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]=0$ and the second factor is finite by Lemma A.5. Thus,
-$|m_{p,i}-\mu_{i}|=\left|\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left(\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]-\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]\right)\right|=\\ \left|\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]\right|\leq\\ \operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{|(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}|}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\,\right]=\\ \operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\,\right]\operatorname{E}\left[\,|(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-m_{p,j})^{2}|\,\right]\leq\\ \frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\cdot\frac{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]^{3/2}}\cdot\\ \operatorname{E}\left[\,\left(\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\right)^{3/2}\,\right]\cdot\sum_{j\in\mathcal{E}}\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|\,\right]\\ =:I\cdot II\cdot III\cdot IV,$
-where we used $\sqrt{x}-\sqrt{y}=\tfrac{x-y}{\sqrt{x}+\sqrt{y}}$, followed by the triangle inequality and $\|\mathbf{X}-\mathbf{m}_{p}\|\geq\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\geq\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|\geq 0$, and finally independence of $(X_{j})_{j\in\mathcal{E}}$ and $(X_{j})_{j\in\{1,\ldots,p\}\setminus\mathcal{E}^{2}(i)}$ because of $M$-independence.
-For the first Term $I$, it holds that $I\leq 1$ using Jensen’s inequality as shown in the proof of Lemma 2.4.
-Regarding Term $II$, note that
-$\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}\,\right])+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2},$
-since $\operatorname{E}\left[\,(\mathbf{X}-\boldsymbol{\mu})^{T}(\boldsymbol{\mu}-\mathbf{m}_{p})\,\right]=0$, and similarly $\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}_{-\mathcal{E}^{2}(i)}\,\right])+\|\boldsymbol{\mu}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|$.
-Using Lemma 2.2, properties of $\mathcal{D}$, and $4M+1\leq\tfrac{1}{3}p$ by $p\geq 14M+8$, we have
-$II\leq\frac{\sqrt{Cp+Cp}}{((p-(4M+1))\sigma_{\min}^{2}+0)^{3/2}}\leq\frac{\sqrt{p}\sqrt{2C}}{(\tfrac{2}{3}p)^{3/2}\,\sigma_{\min}^{3}}=\frac{1}{p}\cdot\frac{\sqrt{2C}}{\sigma_{\min}^{3}\cdot(\tfrac{2}{3})^{3/2}}.$
-Applying Lemma A.5 for $k:=3$ and $n:=2$ to Term $III$, we have $III\leq B_{2}(3,2,\mathcal{D})$.
-For Term $IV$, first note $|\mu_{i}|\leq C$ and $|m_{p,j}|\leq B_{1}+C$ according to Lemma 2.4.
-Thus, applying Jensen’s inequality (note $x\mapsto x^{3/q}$ is concave by $q\geq 3$) and Lemma A.6, we find that for any $j\in\mathcal{E}$ we have
-$\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|\,\right]\leq\operatorname{E}\left[\,(|X_{i}|+C)(|X_{j}|+B_{1}+C)^{2}\,\right]\leq\\ \operatorname{E}\left[\,((|X_{i}|+C)(|X_{j}|+B_{1}+C)^{2})^{q/3}\,\right]^{3/q}\leq B_{4}(C,B_{1}+C,\mathcal{D})^{3/q}.$
-Since $\mathcal{E}$ contains at most $2M+1$ indices, we have shown that term $IV$ is bounded with constant only depending on the $\mathcal{D}$-parameters.
-In summary, Terms $I,III$ and $IV$ are bounded and Term $II$ is $\mathcal{O}(\tfrac{1}{p})$, which yields the Theorem claim for $p\geq 14M+8$.
-Let $p\leq 14M+8$. Then, for all $1\leq i\leq p$, by Lemma 2.2,
-$|m_{p,i}-\mu_{i}|\leq\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\sqrt{Cp}=\frac{1}{p}\sqrt{C}p^{3/2}\leq\frac{1}{p}\sqrt{C}(14M+8)^{3/2}.$
-Taking the maximum of the constants derived in the two cases for $p$ yields the theorem claim. ∎
-See 2.6
-Proof.
-We show
-$p\cdot(m_{p,i}-\mu_{i})\xrightarrow{p\to\infty}\frac{1}{2\bar{\sigma}}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right],$
-which is equivalent to the claim. Similar to the proof of Theorem 2.5, one can see that for $p\geq i+M$
-$p\cdot(m_{p,i}-\mu_{i})=p\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\left(\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]-\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|}\,\right]\right)=\\ p\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]=\\ \frac{1}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\quad\cdot\quad\frac{p\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{3/2}}\quad\cdot\\ \operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{3/2}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\cdot\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\cdot(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right]\\ =:I\cdot II\cdot III.$ (24)
-Starting with the first Term, $I$, we note that it can be rewritten as
-$I=\operatorname{E}\left[\,\left(\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\right)^{-1/2}\,\right]^{-1}.$
-Since by Lemma A.7 we have $\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\to 1$ in probability as $p\to\infty$, also $(\tfrac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]})^{-1/2}\to 1$ by the continuos mapping theorem.
-To show that the random sequence’s expectations also converge, we appeal to a uniform integrability argument [3, Thm 6.2].
-Indeed, by Theorem 2.3 it holds for example that
-$\sup_{p\geq 4(M+1)+1}\operatorname{E}\left[\,\left(\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\right)^{-1}\,\right]\leq\beta(1,C,\mathcal{D})<\infty.$
-[…]
-Regarding Term $II$, recall from the proof of Theorem 2.5 that $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]=\operatorname{tr}(\operatorname{Var}\left[\,\mathbf{X}\,\right])+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}$.
-Further, by Theorem 2.5 it holds that $\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}=\mathcal{O}(\tfrac{1}{p})$ for $p$ large enough.
-Consequently, $\tfrac{1}{p}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\to\bar{\sigma}^{2}$.
-Since $\bar{\sigma}\geq\sigma_{\min}>0$, the second Term converges to $\tfrac{1}{\bar{\sigma}^{2}}$.
-We treat Term $III$ similar to Term $I$, i.e.
-first show convergence in probability of the expectation argument and then appeal to a uniform integrability argument to deduce convergence of the expectations.
-To simplify notation, rewrite $III=\operatorname{E}\left[\,N_{p}/D_{p}\,\right]$, where the numerator $N_{p}$ and denominator $D_{p}$ are given by
-$N_{p}$ $:=-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2},$
-$D_{p}$ $:=\sqrt{\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\cdot\sqrt{\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\left(\sqrt{\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}+\sqrt{\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\right).$
-Starting with the numerator $N_{p}$, note that $m_{p,j}\to\mu_{j}$ by Theorem 2.5.
-Since $\mathcal{E}(i)$ contains finitely many terms, we conclude by the continuous mapping theorem that $N_{p}\to N:=-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}}(X_{j}-\mu_{j})^{2}$ in probability (note that for convergence of a random vector in probability it suffices that the components converge in probability).
-For studying the denominator $D_{p}$, note that for $p\geq i+M$,
-$\frac{\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}=\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}-\frac{\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\xrightarrow{p\to\infty}1-0=1$
-in probability, since the first term converges to $1$ by Lemma A.7, and the numerator of the second term converges to $\sum_{j\in\mathcal{E}(i)}(X_{j}-\mu_{j})^{2}$ as discussed while $\tfrac{1}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\leq\tfrac{1}{p\sigma_{\min}^{2}}\to 0$, implying the second fraction converges to 0 by Slutsky’s theorem.
-Thus, another application of the continuous mapping theorem implies $D_{p}\to 2$ in probability.
-All in all, $N_{p}/D_{p}\to N/2$ in probability by Slutsky’s theorem, where $N$ is as defined previously.
-Next, we show $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$.
-Crucially $q/3>1$, which is required for the uniform integrability argument.
-First, note that similar to the proof of Theorem 2.5,
-$\operatorname{E}\left[\,\left(\frac{|N_{p}|}{D_{p}}\right)^{q/3}\,\right]\leq\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}\operatorname{E}\left[\,\left(\frac{|N_{p}|}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{3}}\right)^{q/3}\,\right]=\\ \frac{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]^{q/2}}{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]^{q/2}}\cdot\operatorname{E}\left[\,|N_{p}|^{q/3}\,\right]\cdot\\ \operatorname{E}\left[\,\left(\frac{\operatorname{E}\left[\,\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}\,\right]}{\|\mathbf{X}_{-\mathcal{E}^{2}(i)}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\right)^{q/2}\,\right]=:III^{\prime}\cdot III^{\prime\prime}\cdot III^{\prime\prime\prime}.$
-We have, using Lemma 2.2, that for $p\geq 4M+2$
-$(III^{\prime})^{2/q}=\frac{\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}\,\right]}+\|\boldsymbol{\mu}-\mathbf{m}_{p}\|^{2}}{\operatorname{tr}{\operatorname{Var}\left[\,\mathbf{X}_{-\mathcal{E}^{2}(i)}\,\right]}+\|\boldsymbol{\mu}-\mathbf{m}_{p,-\mathcal{E}^{2}(i)}\|^{2}}\leq\\ \frac{Cp+Cp}{(p-|\mathcal{E}^{2}(i)|)\sigma_{\min}^{2}+0}\leq\frac{2Cp}{(p-(4M+1))\sigma_{\min}^{2}},$
-which remains bounded as $p\to\infty$ with bound only depending on the $\mathcal{D}$-parameters.
-For the second term, use convexity of $x\mapsto x^{q/3}$, the fact that $|\mu_{i}|\leq C$ and $|m_{p,j}|\leq B_{1}+C$, and Lemma A.6 to find
-$III^{\prime\prime}\leq|\mathcal{E}(i)|^{q/3}\sum_{j\in\mathcal{E}(i)}\frac{1}{|\mathcal{E}(i)|}\operatorname{E}\left[\,|(X_{i}-\mu_{i})(X_{j}-m_{p,j})^{2}|^{q/3}\,\right]\leq\\ |\mathcal{E}(i)|^{(q-3)/3}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,((|X_{1}|+C)(|X_{j}|+B_{1}+C)^{2})^{q/3}\,\right]\leq\\ (2M+1)^{(q-3)/3}(2M+1)B_{4}(C,B_{1}+C,\mathcal{D}).$
-Finally, when $p$ large enough, then by Lemma A.5, we have $III^{\prime\prime\prime}\leq B_{2}(q,2,\mathcal{D})$.
-In summary, $\sup_{p}\operatorname{E}\left[\,|N_{p}/D_{p}|^{q/3}\,\right]<\infty$ when $p$ large enough. By [3, Thm 6.2], we conclude $III\to\operatorname{E}\left[\,N/2\,\right]$ as $p\to\infty$.
-Inserting the obtained limits for $I$, IIII, and $III$ into (24) yields the claim. ∎
+Let $N:=|\{j\in\{1,\ldots,p\}:|j-i|\leq M\}|$ be the number of indices $j$ in the sum. Then, by convexity of $x\mapsto x^{k/3}$, we have that
+$\operatorname{E}\left[\,|\,(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}\,|^{k/3}\,\right]\leq N^{\tfrac{k}{3}-1}\sum_{|j-i|\leq M}\operatorname{E}\left[\,|\,(X_{i}-\mu_{i})(X_{j}-m_{j})^{2}\,|^{k/3}\,\right].$
+Further, by Hölder’s inequality
+$\operatorname{E}\left[\,|\,(X_{i}-\mu_{i})(X_{j}-m_{j})^{2}\,|^{k/3}\,\right]\leq\left(\|X_{i}-\mu_{i}\|_{L^{k}}\cdot\|(X_{j}-m_{j})^{2}\|_{L^{k/2}}\right)^{k/3}\\ =\operatorname{E}\left[\,|X_{i}-\mu_{i}|^{k}\,\right]^{\tfrac{1}{3}}\cdot\operatorname{E}\left[\,|X_{j}-m_{j}|^{k}\,\right]^{\tfrac{2}{3}}.$
+Note $|\mu_{i}|\leq C$ and $|m_{j}|\leq B_{1}+C$ according to Lemma 3.3. Further, for any $A\in\mathbb{R}$ we have that
+$\operatorname{E}\left[\,|X_{i}-A|^{k}\,\right]\leq\operatorname{E}\left[\,(|X_{i}|+|A|)^{k}\,\right]\leq 2^{k-1}(\operatorname{E}\left[\,|X_{i}|^{k}\,\right]+A^{k})\leq 2^{k-1}(C+A^{k}).$
+Taking everything together, we find:
+$\operatorname{E}\left[\,|\,(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}\,|^{k/3}\,\right]\leq N^{\tfrac{k}{3}-1}\cdot N\cdot 2^{k-1}(C+C^{k})^{\tfrac{1}{3}}(B_{1}+2C)^{\tfrac{2}{3}}.$
+Noting $N\leq 2M+1$ yields the claim. ∎
 Lemma A.7.
@@
 If $q<4$,
-$\operatorname{E}\left[\,A_{p,i}^{2}\,\right]=\operatorname{E}\left[\,((X_{i}-m_{p,i})^{2})^{2-q/2}|X_{i}-m_{p,i}|^{q}1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}\,\right]\leq\\ p^{2-q/2}\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq p^{2-q/2}\tilde{C}$.
-Thus, in any case, $\operatorname{Var}\left[\,A_{p,i}\,\right]\leq\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\tilde{C}\cdot\max(1,p^{2-q/2})$.
+$\operatorname{E}\left[\,A_{p,i}^{2}\,\right]=\operatorname{E}\left[\,((X_{i}-m_{p,i})^{2})^{2-q/2}|X_{i}-m_{p,i}|^{q}1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}\,\right]\\ \leq p^{2-q/2}\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq p^{2-q/2}\tilde{C}$.
+Thus, in any case $\operatorname{Var}\left[\,A_{p,i}\,\right]\leq\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\tilde{C}\cdot\max(1,p^{2-q/2})$.
 Let $i\in\mathbb{N}$.
 By $M$-dependence, $\operatorname{Cov}(A_{p,i},A_{p,j})=0$ for all but $2M+1$ indices $j$.
-This fact together with $|\operatorname{Cov}(X,Y)|\leq\sqrt{\operatorname{Var}\left[\,X\,\right]\operatorname{Var}\left[\,Y\,\right]}$ gives
+This fact together with $|\operatorname{Cov}(X,Y)|\leq\sqrt{\operatorname{Var}\left[\,X\,\right]\operatorname{Var}\left[\,Y\,\right]}$ gives:
 $\left|\sum_{j=1}^{p}\operatorname{Cov}(A_{p,i},A_{p,j})\right|\leq(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2}).$
-Thus,
+Thus:
 $\operatorname{E}\left[\,\left(\sum_{i=1}^{p}A_{p,i}-a_{p}\right)^{2}\,\right]=\operatorname{Var}\left[\,\sum_{i=1}^{p}A_{p,i}\,\right]\leq p\cdot(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2})$
@@
 $\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right)\leq\frac{p\cdot(2M+1)\cdot\tilde{C}\cdot\max(1,p^{2-q/2})}{t^{2}a_{p}^{2}}.$
-Note
+Note that
 […]
@@
 $\operatorname{P}\left(\,\left|\sum_{i=1}^{p}A_{p,i}-a_{p}\right|>a_{p}t\,\right)\leq\frac{1}{t^{2}}\cdot\frac{4(2M+1)\tilde{C}}{\sigma_{\min}^{4}}\cdot\max\left(\frac{1}{p},p^{1-q/2}\right)\xrightarrow{p\to\infty}0.$ (30)
-Combining (30) with (29) in (28), we have shown
+Combining (30) with (29) in (28), we have shown that
 $\frac{\|\mathbf{X}-\mathbf{m}_{p}\|^{2}}{a_{p}}\xrightarrow{P}1$ (31)
-in probability as $p\to\infty$. Further, note
+in probability as $p\to\infty$. Further, note that
 $\frac{\sum_{i=1}^{p}\operatorname{E}\left[\,(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}>p\}}\,\right]}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\leq\frac{p^{2-q/2}}{p\,\sigma_{\min}^{2}}\left(\tilde{C}+\frac{\tilde{C}}{q/2-1}\right)\xrightarrow{p\to\infty}0,$
-which implies
+which implies that
 $\frac{a_{p}}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}=1-\frac{\sum_{i=1}^{p}\operatorname{E}\left[\,(X_{i}-m_{p,i})^{2}\cdot 1_{\{(X_{i}-m_{p,i})^{2}>p\}}\,\right]}{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}\xrightarrow{p\to\infty}1.$
 Combining this with (31) yields the claim by Slutsky’s theorem. ∎
-A.2 Further analysis of the simulation results
+Appendix B Further analysis of the simulation results
 Linear Regression slopes for the simulation curves can be found in Table 2. The Pareto(3.1) results are analyzed in more detail in Figure 2.